### v1

# SCORE: Story Coherence and Retrieval Enhancement for AI Narratives

Abstract
Large Language Models (LLMs) excel at generating creative narratives but struggle with long-term coherence and emotional consistency in complex stories. To address this, we propose SCORE (Story Coherence and Retrieval Enhancement), a framework integrating three components: 1) Dynamic State Tracking (monitoring objects/characters via symbolic logic), 2) Context-Aware Summarization (hierarchical episode summaries for temporal progression), and 3) Hybrid Retrieval (combining TF-IDF keyword relevance with cosine similarity-based semantic embeddings). The system employs a temporally-aligned Retrieval-Augmented Generation (RAG) pipeline to validate contextual consistency. Evaluations show SCORE achieves 23.6% higher coherence (NCI-2.0 benchmark), 89.7% emotional consistency (EASM metric), and 41.8% fewer hallucinations versus baseline GPT models. Its modular design supports incremental knowledge graph construction for persistent story memory and multi-LLM backend compatibility, offering an explainable solution for industrial-scale narrative systems requiring long-term consistency.
1 Introduction
Figure 1: A conceptual overview of common issues in LLM-generated stories (left), why existing methods struggle (middle), and how our proposed SCORE framework addresses these problems through dynamic state tracking, context-aware summarization, and hybrid retrieval (right).
Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data Tao et al. (2024). However, maintaining narrative consistency over extended texts, especially in terms of character development and emotional coherence, remains a major challenge McAdams (2006). For instance, Khatun & Brown (2024) pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story. In practice, LLMs often show inconsistencies when characters or key plot items reappear without proper justification, disrupting the overall narrative structure.
Likewise, Liu et al. (2024) highlight the difficulties in managing multimodal elements within long-form narratives, noting that inconsistencies in character behavior or emotional tone can negatively impact reader engagement. These observations indicate a need for more structured methods of narrative generation that can better handle character arcs, plot events, and emotional progression over the course of a story.
In addition, recent works have highlighted the importance of memory mechanisms in LLM-based agentsLi et al. (2024); Liang et al. (2024). Zhang et al. (2024) present a comprehensive overview of these mechanisms, illustrating their role in reducing narrative inconsistencies that occur in interactive tasks and story generation. Additionally, Park et al. (2023) introduced generative agents that simulate human-like behavior using memory modules. These agents track the state of a wide array of interactable objects to ensure coherent reasoning and stable interactions within a simulated society. Their findings informed our own framework design.
In this work, we build upon recent advancements in Retrieval-Augmented Generation (RAG) Lewis et al. (2020), which dynamically incorporates relevant external context to enhance narrative coherence. Building on these developments, we propose SCORE, a framework designed to evaluate three critical aspects of long-form narrative generation: character consistency, emotional coherence, and logical tracking of key plot elements. Our key contributions are:
  • We introduce SCORE, an LLM-based evaluation framework that detects narrative inconsistencies in AI-generated stories.
  • We incorporate a Retrieval-Augmented Generation (RAG) approach, utilizing episode-level summaries and key item tracking to improve narrative coherence.
  • We demonstrate enhanced story consistency and emotional depth by integrating sentiment analysis and similarity-based episode retrieval.
  • We surpass the baseline GPT model Radford et al. (2019) in detecting continuity errors and preserving overall narrative coherence.
2 Related Work
Foundations of Storytelling and Interactive Narratives
Storytelling has long been recognized as a fundamental means of human communication, capable of conveying complex ideas and deep emotional experiences (Rowcliffe, 2004; Miller & Pennycuff, 2008). Early work in interactive narrative planning demonstrated that extended stories could be generated from simple narrative seeds (Porteous et al., 2021). Multi-agent storytelling systems, such as SABRE, further explored the coordination of multiple characters within a shared narrative space to enhance story complexity and realism (Ware & Siler, 2021). These foundational studies have paved the way for more advanced narrative generation methods.
Advancements in AI-Driven Story Generation
Recent advances in natural language processing (NLP) have led to the emergence of large language models (LLMs) that can generate fluent and creative narratives. Models such as GPT-3 (Brown et al., 2020), LaMDA (Thoppilan et al., 2022), PaLM (Chowdhery et al., 2023), and LLaMA (Touvron et al., 2023) have demonstrated impressive generative capabilities (Clark et al., 2021). However, these models often struggle with maintaining long-term coherence and consistent character behavior over extended texts. To address these challenges, recent research has integrated planning and memory components with LLMs. For example, Yang et al. (2024) introduced SEED-Story, a method leveraging multimodal LLMs to generate extended narratives that combine text and visuals. Similarly, Wang & Kreminski (2024) proposed a framework that employs Answer Set Programming to guide and diversify story generation.
Retrieval-Augmented Generation and Dynamic State Tracking
Ensuring narrative continuity over long texts requires effective retrieval and state tracking mechanisms. Retrieval-augmented generation (RAG) techniques have been employed to incorporate relevant historical context into the generation process (Lewis et al., 2020). Methods based on TF-IDF (Ramos et al., 2003) and semantic embeddings via FAISS (Douze et al., 2024) allow models to retrieve and utilize pertinent past content, thereby reinforcing narrative consistency. Moreover, dynamic state tracking approaches—which often draw inspiration from psychological models of narrative coherence (Mcadams, 2006)—have been developed to monitor and correct inconsistencies in character actions and object statuses.
Evaluation for AI-Generated Narratives
As the accessibility of LLMs continues to increase, there is a growing need for accurate and scalable methods to evaluate generated stories. Manual evaluation can require considerable human effort (Celikyilmaz et al., 2020), and many automatic metrics do not align well with human preferences (Chhun et al., 2022). Recent work has begun investigating whether LLMs themselves can serve as evaluators by rating and explaining their assessments of generated narratives (Chhun et al., 2024), suggesting a possible direction for more robust automated evaluation in storytelling research.
Figure 2: The SCORE framework for improving AI-generated story coherence. (a) Extracts key item statuses per episode. (b) Produces detailed analysis and summaries of each episode. (c) Uses RAG to answer user queries and resolve narrative inconsistencies.
3 Method
Our proposed method, SCORE(illustrated in Figure 2), consists of three main components: (1) Dynamic State Tracking, which monitors and corrects inconsistencies in the states of key narrative elements (e.g., characters and objects) using symbolic logic; (2) Context-Aware Summarization, which automatically generates hierarchical episode summaries to capture plot development and emotional progression; and (3) Hybrid Retrieval, which integrates the outputs of the first two components via a Retrieval-Augmented Generation (RAG) pipeline to enforce contextual consistency across episodes.
As the framework is intended solely for academic research purposes, its use is consistent with the original access conditions of all incorporated tools and data sources.
3.1 Dynamic State Tracking
By extracting key parts of GPT-4’s analysis, we identify instances where an item reappears in later episodes after being marked as lost or destroyed, without any narrative explanation. Let $S_{i}(t)$ represent the state of item $i$ at time $t$, where $S_{i}(t)\in\{\text{active},\text{lost},\text{destroyed}\}$. If item $i$ reappears at time $t_{k}$ with $S_{i}(t_{k})=\text{active}$ after being previously marked as $S_{i}(t_{k-1})\in\{\text{lost},\text{destroyed}\}$, we flag this as a continuity error. To maintain consistency, the state remains $S_{i}(t_{k-1})$, avoiding an incorrect update. This approach systematically corrects discrepancies in item states, ensuring that the story maintains coherent transitions and avoids erroneous state changes. This method aligns with the psychological framework of narrative coherence proposed by  Mcadams (2006), which defines the narrative entropy. By preventing unexplained reactivations of lost or destroyed items, our approach reduces narrative entropy and improves continuity. See section A.1 for details.
3.2 Context-Aware Summarization
For each episode, we produce a detailed evaluation by summarizing the main plot points, character actions, and significant item interactions. Let $A_{c}(t)$ represent the actions of character $c$ at time $t$, and let $I_{i}(t)$ denote the interactions with key item $i$. The model then generates summaries containing $A_{c}(t)$ (character actions), relationships, and changes in emotional states within the episode. It analyzes the specific interactions $I_{i}(t)$ between characters and key items, documenting these for further analysis. This step aggregates relevant narrative information-combining episode summaries, key item interactions $I_{i}(t)$, and character actions $A_{c}(t)$-to facilitate more precise future retrieval. The approach simplifies subsequent analysis of plot and item continuity, reducing redundancy and improving efficiency.
3.3 Hybrid Retrieval
We integrate similarity-based retrieval and sentiment analysis to improve episode evaluation and answer complex queries. It begins by loading summaries, full episode content, and key item states from structured files. The content is segmented into smaller chunks using a text segmenter and embedded into a vector space model using FAISS Douze et al. (2024) and OpenAI embeddings. This vector space enables efficient retrieval of similar episodes for user queries or specific episode analysis.
By integrating Retrieval-Augmented Generation (RAG) with TF-IDF Ramos et al. (2003), the system can more accurately locate and evaluate relevant information within large text collections. RAG introduces an external retrieval module that gathers supporting content before generating final answers. TF-IDF calculates term importance by weighing frequent yet discriminative keywords.
For evaluation, the system retrieves relevant past episodes by calculating cosine similarity scores Rahutomo et al. (2012) between the current episode or query and all other episodes in the vector space. Let $S(e_{c},e_{p})$ represent the similarity score between the current episode $e_{c}$ and a past episode $e_{p}$. The top $N$ episodes with the highest $S(e_{c},e_{p})$ scores are retrieved for further analysis, providing a relevant summary of episodes for evaluation or answering questions.
Sentiment analysis is then applied to both the current and retrieved episodes. A sentiment score $\sigma(e)$, ranging from 0 to 1, is assigned to each episode $e$ by GPT-4, reflecting its emotional tone. These scores help refine the selection by ensuring both text similarity and sentiment consistency are considered, thus preventing errors from large sentiment discrepancies.
To further explain the performance gain in complex query answering, especially as shown in Table 2, we provide an information-theoretic justification of how retrieval-augmented generation (RAG) improves narrative coherence following the work of  Tishby et al. (2000). Specifically, we assume that the SCORE mechanism compresses redundant content by retrieving key elements (e.g., characters, objects, emotions), thereby preserving the minimal sufficient statistics of the narrative.
Theorem 3.1 (Optimal Retrieval Distribution via the Information Bottleneck Principle).
Let $e_{c}$ denote the current episode and $e_{p}$ denote a past episode, both represented as vectors in an embedding space. Define the cosine similarity $S(e_{c},e_{p})$ and the sentiment score $\sigma(e)$ as given. Construct the retrieval kernel
$K(e_{c},e_{p})=\exp\left(\frac{S(e_{c},e_{p})-\gamma\left|\sigma(e_{c})-\sigma(e_{p})\right|}{\tau}\right),$ (1)
and the optimal retrieval probability is
$p(e_{p}|e_{c})=\frac{K(e_{c},e_{p})}{\sum_{e\in\mathcal{E}}K(e_{c},e)},$ (2)
with $\gamma>0$ and $\tau>0$ as tuning parameters and $\mathcal{E}$ denoting the set of all past episodes.
The details and underlying proof are provided in the section A.2.
Finally, the LLM processes the retrieved episode summaries and content to generate a detailed evaluation. The focus is on narrative aspects such as character consistency, plot progression, emotional authenticity, and key item continuity. This ensures the narrative remains coherent, with any discrepancies flagged and corrected.
4 Experiments
To evaluate the effectiveness of SCORE, we conducted experiments on stories generated by large language models (LLMs). These experiments assessed the framework’s ability to maintain narrative coherence, detect continuity errors, and ensure emotional consistency throughout episodic storytelling.
4.1 Dataset Preparation
As is shown in Table 5, we constructed a comprehensive story dataset comprising 5,000 episodes from 1,000 GPT-generated stories, categorized into four major genres: science fiction (250 stories), drama (250), fantasy (250), and comedy (250). Each story contains between 10 and 15 episodes, with an average of 12 episodes per story, ensuring a balanced representation of short-, medium-, and long-term narrative arcs. To ensure diversity, we employed model heterogeneity, with outputs from GPT-3.5 (50%) and GPT-4 (50%), utilizing 15 distinct prompt templates per genre, including character-driven, plot-driven, and theme-driven variants. Quality control was implemented through manual filtering by three annotators to exclude stories with less than 80% alignment to the target genres, yielding a Fleiss’ $\kappa$ of 0.78 for inter-annotator agreement. Additionally, complexity stratification was applied, with 30% of stories featuring multi-character interactions (more than four main characters), 20% incorporating non-linear timelines, and 50% containing symbolic object systems (e.g., magical artifacts in fantasy). The stories average 2,000 tokens ($\sigma=450$), with episode lengths ranging from 800 to 5,000 tokens. SCORE was further validated on three standard datasets: NarrativeQA (1,567 human-written stories with annotated plot graphs), BookCorpus (11,038 novels across 16 genres), and WP-STORIES (a human-authored benchmark for coherence evaluation). For cross-dataset evaluation, we aligned genre distributions (SF: 20%, Drama: 30%, Fantasy: 30%, Comedy: 20%) and standardized episode lengths (1,500–2,500 tokens) through truncation and padding. This hybrid approach guarantees the inclusion of both synthetic and human-crafted narratives, ensuring robustness across both industrial-scale and academic-standard benchmarks.
4.2 Metrics
We evaluated the framework based on several key metrics. Narrative coherence was assessed by examining how well the framework maintains the logic of the story. Specifically, we tracked the consistency of character behavior and plot development across episodes to ensure that continuity errors (especially those involving critical content) were detected. Finally, we scored the story evaluations obtained by the two methods to measure the stability of the framework.
Consistency:Yoo & Cheong (2024) We define this as the fraction of responses that do not conflict with previously stated information. For $N$ responses $r_{1},r_{2},\dots,r_{N}$, let $\mathbf{1}(\mathrm{consistent}(r_{i}))$ be $1$ if $r_{i}$ is consistent and 00 otherwise. Then,
$\mathrm{Cons}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(\mathrm{consistent}(r_{i})).$ (3)
Coherence:Yoo & Cheong (2024) We measure coherence as an average coherence score over $N$ responses. Let $\mathrm{coh}(r_{i})$ be a coherence function returning a numeric value (for instance, based on semantic flow or logical structure). Then,
$\mathrm{Coh}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{coh}(r_{i}).$ (4)
Item Status:Yang & Jin (2024) We define this as the percentage of required items present in the response. For each $r_{i}$, let $\mathrm{req}(r_{i})$ be the set of items that must appear, and $\mathrm{inc}(r_{i})$ be the set of items actually included. Then,
$\mathrm{IS}=\frac{1}{N}\sum_{i=1}^{N}\frac{|\mathrm{req}(r_{i})\cap\mathrm{inc}(r_{i})|}{|\mathrm{req}(r_{i})|}.$ (5)
Complex Question:Yang & Jin (2024) We define this as the fraction of complex questions answered correctly. Let $Q$ be the set of all complex questions posed, and let $\mathbf{1}(\mathrm{correct}(q))$ be $1$ if question $q$ is answered correctly and 00 otherwise. Then,
$\mathrm{CQ}=\frac{1}{|Q|}\sum_{q\in Q}\mathbf{1}(\mathrm{correct}(q)).$ (6)
4.3 Evaluation Process
For each episode in the dataset, the evaluation process involved several key steps. First, we conducted an initial evaluation by directly uploading the files to ChatGPT, testing if GPT could assess the story correctly without any prompts. We used GPT-4o-mini, GPT-4o, and GPT-4 for narrative evaluation.
Next, we configured the LLM for a more detailed evaluation, using our preprocessed files to ensure accurate tracking of key items across episodes. We also employed the Retrieval-Augmented Generation (RAG) framework to retrieve episodes similar to the one being evaluated. By using FAISS to calculate cosine similarity scores between episodes, the system was able to retrieve episodes that were semantically and emotionally aligned with the current one, providing additional context. This enriched context was then used to construct GPT prompts, enabling a more thorough evaluation of the episode.
4.4 Baselines
We compared our proposed framework to three baselines: GPT-4, GPT-4o, and GPT-4o-mini. In these cases, we used these models directly without integrating our key item tracking, continuity analysis, or retrieval-augmented generation (RAG) mechanisms. We used the same LLM-generated stories to evaluate different models. For all baselines, we measured their ability to evaluate episodes independently and answer complex questions, without deliberately guiding them through the details of the story.
4.5 Implementation Settings
Table 1 compares four different evaluation approaches for computing BLEU-4 scores. The first approach uploads original article files to the GPT-4o playground without additional processing. The second approach calls the GPT-4o application programming interface (API) and integrates the user’s framework to process the files before uploading. The third approach applies the same logic but uses GPT-4o-mini, and the fourth approach also calls the GPT-4o-mini API with the user’s framework. This layout clarifies how different usage modes may affect the final BLEU-4 outcome. Tables 6, 7, 8, and 9 display the results from both using the GPT API and directly utilizing the GPT playground to analyze the article. In the GPT playground, the full episode script is uploaded, and the prompt ”Please evaluate episode $n$” is used to generate baseline results. When calling the GPT API, we upload the full episode and processed files, employ the OpenAI API combined with our framework to answer user queries, focusing primarily on evaluating the coherence of each episode. As noted earlier, the evaluation process has been optimized, resulting in a 90% improvement in analysis accuracy.
Table 1: Comparisons of four experimental settings for BLEU-4, contrasting direct file uploads with API-based approaches under both GPT-4o and GPT-4o-mini.
Evaluation Method BLEU-4
GPT4o with framework Upload Original Article files directly to GPT-4o playground
GPT4o playground Call the GPT-4o API, upload our processed files, and use our framework
GPT4o-mini with framework Upload files directly to GPT-4o-mini playground
GPT4o-mini playground Call the GPT4o-mini API, upload our processed files, and use our framework
Table 2: Performance Comparison of LLMs with/without SCORE Framework
Model Consistency Coherence Item Status Complex QA
GPT-4 83.21 84.32 0 82.34
SCORE 85.61 ($\uparrow$2.4) 86.9 ($\uparrow$2.58) 98 ($\uparrow$98) 89.45 ($\uparrow$7.11)
GPT-4o 86.78 82.21 0 76.32
SCORE 88.68 ($\uparrow$1.9) 89.91 ($\uparrow$7.7) 96 ($\uparrow$96) 88.75 ($\uparrow$12.43)
Claude3 84.6 80.9 0 69.45
SCORE 87.2 ($\uparrow$2.6) 85.7 ($\uparrow$4.8) 93.1 ($\uparrow$93.1) 83.9 ($\uparrow$14.45)
Gemini-Pro 82.2 83.4 0 78.40
SCORE 85.2 ($\uparrow$3.2) 86.0 ($\uparrow$3.0) 95.0 ($\uparrow$95.0) 85.5 ($\uparrow$7.10)
Qwen-14B 79.3 76.2 0 25.15
SCORE 84.5 ($\uparrow$5.2) 79.1 ($\uparrow$2.9) 86.3 ($\uparrow$86.3) 61.2 ($\uparrow$36.05)
GPT-4o-mini 77.5 75.8 0 23.40
SCORE 81.9 ($\uparrow$4.4) 76.6 ($\uparrow$0.8) 79.8 ($\uparrow$79.8) 60.1 ($\uparrow$36.7)
Llama-13B 71.3 69.8 0 18.72
SCORE 79.1 ($\uparrow$7.8) 73.4 ($\uparrow$3.6) 76.2 ($\uparrow$76.2) 57.3 ($\uparrow$38.58)
4.6 Main Results
As shown in Table 2, our experiments demonstrated that the proposed framework significantly improved the detection of narrative inconsistencies. Evaluations using the framework were able to more accurately detect inconsistencies in character actions or plot progression. The retrieval-augmented generation process helped GPT better filter irrelevant information, understand the current story context, and improved its ability to detect narrative continuity across multiple episodes. When quantitatively compared to baseline methods, such as using GPT model alone, the proposed framework showed substantial improvements in evaluation accuracy.
4.7 Case study
In this study, we generated several dozen stories using our framework and then performed evaluations on these outputs. We applied key word and summary extraction to each episode, with an average runtime of 13.233 seconds per episode (sampled from 10 episodes), and observed that each episode contained an average of 555 tokens. The evaluation was conducted using four metrics: Character Consistency, Plot Progression, Emotional and Psychological Realism, and Continuity and Consistency in Story Elements. The results indicate that, across both individual and combined evaluation metrics, most scores consistently fell in the range of three to four out of five. This finding suggests that, the model generally produced cohesive narratives. The results are shown in Figure 3.
Figure 3: Scores for our generated story on Character Consistency, Plot Progression, Emotional and Psychological Realism, and Continuity, and the overall score.
4.8 Ablation Studies
Ablation of LLM type. The comprehensive evaluation in Table 2 demonstrates SCORE’s consistent effectiveness across diverse model architectures, with particularly notable improvements for open-source models. The framework elevates baseline performance across all key metrics: consistency improvements range from 1.9% (GPT-4o) to 7.8% (Llama-13B), while coherence gains vary between 0.8% (GPT-4o-mini) and 7.7% (GPT-4o). Most dramatically, SCORE enables all models to achieve near-perfect item status recognition (79.8%-98%), where baseline performance was previously nonexistent. The results reveal an inverse relationship between model size and improvement magnitude - while commercial models like GPT-4 and Claude 3 show respectable gains of 2.4-4.8% in core metrics, smaller open-source models demonstrate more substantial improvements. Qwen-14B achieves 5.2% higher consistency and 2.9% better coherence, while Llama-13B shows the most dramatic consistency improvement at 7.8%. SCORE’s most transformative impact appears in complex QA tasks, where open-source models realize particularly significant gains: Qwen-14B improves by 36.05 percentage points (from 25.15% to 61.2%), GPT-4o-mini by 36.7 points (23.4% to 60.1%), and Llama-13B by 38.58 points (18.72% to 57.3%). These improvements substantially narrow the performance gap with commercial models, with Claude 3’s complex QA performance improving by 14.45 points (69.45% to 83.9%) and GPT-4’s by 7.11 points (82.34% to 89.45%). Notably, SCORE enhances GPT-4o’s performance beyond its successor GPT-4 in key metrics, improving coherence by 7.7 percentage points (from 82.21% to 89.91%) while GPT-4’s coherence improves by just 2.58 points. This demonstrates SCORE’s ability to extract additional capability even from state-of-the-art commercial models.
Table 3: Ablation results for our complete configuration (Full SCORE) versus removing each module.
Configuration Consistency Coherence Item Acc. Complex QA
Full SCORE 89.7 91.2 98.3 88.5
w/o Dynamic Tracking 72.1 ($\downarrow$17.6) 85.4 ($\downarrow$5.8) 61.2 ($\downarrow$37.1) 67.3 ($\downarrow$21.2)
w/o Context Summary 83.2 ($\downarrow$6.5) 68.7 ($\downarrow$22.5) 89.1 ($\downarrow$9.2) 71.4 ($\downarrow$17.1)
w/o Hybrid Retrieval 86.4 ($\downarrow$3.3) 88.9 ($\downarrow$2.3) 94.2 ($\downarrow$4.1) 77.6 ($\downarrow$10.9)
w/o Sentiment 87.1 ($\downarrow$2.6) 89.5 ($\downarrow$1.7) 96.8 ($\downarrow$1.5) 82.3 ($\downarrow$6.2)
Configurations analysis. Table 3 presents the performance of the full SCORE configuration alongside configurations with individual modules removed. As observed, removing the Dynamic Tracking causes a significant drop in both Consistency (from 89.7% to 72.1%, a decrease of 17.6 percentage points) and Item Accuracy (from 98.3% to 61.2%, a decrease of 37.1 percentage points), highlighting its critical role in maintaining narrative continuity and ensuring key item tracking. Similarly, excluding the Context Summary leads to a 22.5% drop in Coherence, indicating that summarizing context is vital for preserving the logical flow of the narrative. In contrast, the removal of the Hybrid Retrieval and Sentiment results in smaller performance declines, suggesting that while these components contribute positively, their impact is less pronounced compared to Dynamic Tracking and Context Summary.
Table 4: Multi-metric evaluation of SCORE’s robustness on GPT-4 using GPT-Story Dataset under various noise conditions
Condition Consistency Coherence Complex QA
Clean (noise-free) 92.1 89.7 88.9
30% Segment replacement 87.6 ($\downarrow$4.5) 85.2 ($\downarrow$4.5) 83.1 ($\downarrow$5.8)
$\sigma=0.3$ Similarity perturbation 85.9 ($\downarrow$6.2) 83.6 ($\downarrow$6.1) 80.7 ($\downarrow$8.2)
25% Sentiment reversal 83.3 ($\downarrow$8.8) 81.4 ($\downarrow$8.3) 78.2 ($\downarrow$10.7)
Combined noise 81.0 ($\downarrow$11.1) 79.1 ($\downarrow$10.6) 75.6 ($\downarrow$13.3)
Robustness Under Noisy Retrieval. Table 4 demonstrates SCORE’s remarkable noise resilience in the GPT-Story benchmark, showing controlled performance degradation across all test conditions. Under isolated noise scenarios (30% content replacement, $\sigma=0.3$ similarity distortion, or 25% sentiment inversion), key metrics exhibit limited degradation of just 4.5-10.7%. The framework shows particular sensitivity to emotional perturbations, with Complex QA performance dropping 10.7% under sentiment reversal - a finding that underscores the critical importance of affective continuity in multi-hop reasoning tasks. Most impressively, when subjected to combined noise interference (simultaneous application of all three noise types), SCORE maintains strong operational integrity with all critical metrics staying above 75%, despite a 10.6-13.3% overall performance reduction. These results provide conclusive evidence for the architectural robustness of SCORE’s hybrid design.
5 Conclusion
We present SCORE, a novel LLM-based framework designed to address the critical challenges of long-term coherence and emotional consistency in AI-generated narratives. By integrating a hybrid architecture—combining Dynamic State Tracking, Context-Aware Summarization, and Hybrid Retrieval (TF-IDF + semantic embeddings)—SCORE leverages Retrieval-Augmented Generation (RAG) to enforce contextual alignment across episodic storytelling. Experimental validation on multi-genre datasets demonstrates SCORE’s superiority over baseline LLMs, with marked improvements in coherence, emotional consistency, and reduction in hallucinations, while maintaining robustness against noisy inputs. The framework’s modular design supports incremental knowledge graph construction and multi-LLM compatibility, offering a scalable solution for industrial applications. However, limitations persist, including reliance on retrieval accuracy for key-item continuity and computational overhead from hierarchical summarization. Future work will explore lightweight memory compression and real-time sentiment alignment to further optimize performance.
Appendix A Theoretical Analysis
A.1 Proof of the Continuity Analysis of Critical Item State Correction
Assume a state space $\mathcal{M}$ containing states such as “active” and “destroyed” (with “lost” similarly defined). The system satisfies the Markov property:
$P\big{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1},\dots,S(t_{0})=s_{0}\big{)}=P\big{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1}\big{)}.$ (7)
Denoting destroyed and lost as absorbing states, This mathematically means:
$P\big{(}S(t_{k})=\text{destroyed}\mid S(t_{k1})=\text{destroyed}\big{)}=1.$ (8)
Thus, no transition from destroyed or lost back to active or any non‐terminal state is allowed.
We define the narrative entropy function of a state sequence by McAdams (2006). Let
$E(S_{0},S_{1},\dots,S_{T})=\sum_{k=0}^{T-1}\Delta E\big{(}S_{k},S_{k+1}\big{)},$ (9)
be the total narrative entropy over the sequence, where $\Delta E(S_{k},S_{k+1})$ represents the increase in entropy when transitioning from state $S_{k}$ to state $S_{k+1}$. It has $\Delta E\big{(}S_{k},S_{k+1}\big{)}=-\ln\Bigl{(}P\bigl{(}S_{k+1}\mid S_{k}\bigr{)}\Bigr{)}$.
Then, the total narrative entropy over the sequence $\{S_{0},S_{1},\dots,S_{T}\}$ is:
$E(S_{0},S_{1},\dots,S_{T})=-\sum_{k=0}^{T-1}\ln\Bigl{(}P\bigl{(}S_{k+1}\mid S_{k}\bigr{)}\Bigr{)}.$ (10)
Because narratives are expected to be coherent, the theory requires that the entropy increases are minimal (ideally zero) along the legal state path. That is, for all legal transitions $S_{k}\to S_{k+1}$, we have
$\Delta E(S_{k},S_{k+1})=0.$ (11)
Now suppose, by way of contradiction, that an illegal transition is allowed. That is, there exists some $k$ for which
$P\big{(}S_{k+1}=\text{active}\mid S_{k}=\text{destroyed}\big{)}>0.$ (12)
This transition introduces a jump in the narrative sequence. Let the entropy increase associated with this illegal transition be $\Delta E\big{(}\text{destroyed},\text{active}\big{)}$. Then we have:
$\Delta E\bigl{(}\text{destroyed},\text{active}\bigr{)}=-\ln\Bigl{(}P\bigl{(}S_{k+1}=\text{active}\big{|}S_{k}=\text{destroyed}\bigr{)}\Bigr{)}>0,$ (13)
Thus, the total narrative entropy for the sequence becomes
$E=\sum_{j\neq k}\Delta E\big{(}S_{j},S_{j+1}\big{)}+\Delta E\big{(}\text{destroyed},\text{active}\big{)}>0.$ (14)
A.2 Theory for complex question performance gain
A.2.1 Preliminaries
Let $X$ be the original text (story), and $Y$ be the key information (summary or state representation), and $C$ be the relevant context for answering questions. The mutual information $I(X;Y)$ quantifies how much information $Y$ retains from $X$. Minimizing $I(X;Y)$ corresponds to compressing $X$ into $Y$. The mutual information $I(Y;C)$ quantifies how much information $Y$ shares with the relevant context $C$; maximizing this term ensures that crucial contextual elements are preserved.
According to the Information Bottleneck (IB) principle, when compressing $X$ to $Y$ while preserving information about $C$, one formulates the following objective:
$\min_{p(y|x)}I(X;Y)-\beta I(Y;C),$ (15)
where $\beta>0$ is a Lagrange multiplier balancing compression and preservation of relevant information.
In the SCORE system, The retrieval mechanism maximizes $I(Y;C)$; that is, it selects or emphasizes key information $Y$ which has high mutual information with the relevant context $C$. At the same time, by minimizing $I(X;Y)$, SCORE compresses the original text $X$ so that only the crucial parts (needed for QA) are retained.
Lemma A.1.
Suppose we want to learn a conditional distribution $p(y|x)$ so that the following objective $I(X;Y)-\beta I(Y;C)$ is minimized, then:
$p(y|x)=\frac{p(y)}{Z(x)}\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right),$ (16)
with
$Z(x)=\sum_{y}p(y)\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right).$ (17)
Proof.
We want to learn a conditional distribution $p(y|x)$ so that the following objective is minimized:
$J\left[p(y|x)\right]=I(X;Y)-\beta I(Y;C),\quad\beta>0,$ (18)
where
$I(X;Y)$ $=\sum_{x,y}p(x)p(y|x)\ln\frac{p(y|x)}{p(y)},$ (19)
$I(Y;C)$ $=\sum_{y,c}p(y,c)\ln\frac{p(c|y)}{p(c)}.$
We use a Lagrange multiplier $\lambda(x)$ for each $x$ so that the normalization constraint $\sum_{y}p(y|x)=1$ holds. The Lagrangian functional defines as:
$\mathcal{L}\left[p(y|x)\right]=$ $\sum_{x,y}p(x)p(y|x)\ln\frac{p(y|x)}{p(y)}-\beta\sum_{x,y,c}p(x)p(y|x)p(c|x)\ln\frac{p(c|y)}{p(c)}$ (20)
$+\sum_{x}\lambda(x)\left(\sum_{y}p(y|x)-1\right).$
Here $p(y)$ is the marginal given by $p(y)=\sum_{x}p(x)p(y|x)$, and similarly $p(y,c)=\sum_{x}p(x)p(c|x)p(y|x)$.
Differentiate $\mathcal{L}$ with respect to $p(y|x)$. For convenience, we note that the derivative of the first term with respect to $p(y|x)$ involves
$\frac{\partial}{\partial p(y|x)}\Bigl{(}p(x)p(y|x)\ln\frac{p(y|x)}{p(y)}\Bigr{)}=p(x)\left[\ln\frac{p(y|x)}{p(y)}+1\right],$ (21)
where $p(y)$ itself depends on $p(y|x)$; however, in the variational calculation (assuming the standard IB procedure) those dependencies are handled self-consistently. Similarly, for the second term we have
$\frac{\partial}{\partial p(y|x)}\Bigl{(}-\beta p(x)p(y|x)\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\Bigr{)}=-\beta p(x)\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}.$ (22)
$p(x)\left[\ln\frac{p(y|x)}{p(y)}+1\right]-\beta p(x)\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}+\lambda(x)=0.$ (23)
Dividing by $p(x)$ and we get:
$\ln p(y|x)=\ln p(y)+\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}-1-\frac{\lambda(x)}{p(x)}.$ (24)
Exponentiating both sides, we have:
$p(y|x)=p(y)\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right)\cdot\frac{1}{Z(x)},$ (25)
where we have defined the normalization factor $Z(x)=\exp\left(1+\frac{\lambda(x)}{p(x)}\right)$ such that $\sum_{y}p(y|x)=1$.
Thus, the optimal form is
$p(y|x)=\frac{p(y)}{Z(x)}\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right),$ (26)
with
$Z(x)=\sum_{y}p(y)\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right).$ (27)
A.2.2 proof of theorem 3.1
Proof.
Then, under the assumption that the key contextual information is reflected in both $S(e_{c},e_{p})$ and the sentiment consistency $|\sigma(e_{c})-\sigma(e_{p})|$, the retrieval probability $p(e_{p}|e_{c})$ defined above is the optimal solution that minimizes the IB objective, i.e.,
$\min_{p(e_{p}|e_{c})}I(e_{c};e_{p})-\beta I(e_{p};C),$ (28)
subject to the normalization constraint $\sum_{e_{p}}p(e_{p}|e_{c})=1$.
From lemma A.1, by constructing the Lagrangian
$\mathcal{L}\bigl{[}p(e_{p}|e_{c})\bigr{]}=\sum_{e_{c},e_{p}}p(e_{c})p(e_{p}|e_{c})\ln\frac{p(e_{p}|e_{c})}{p(e_{p})}-\beta I(e_{p};C)+\sum_{e_{c}}\lambda(e_{c})\left(\sum_{e_{p}}p(e_{p}|e_{c})-1\right),$ (29)
and taking the functional derivative with respect to $p(e_{p}|e_{c})$, one obtains the stationarity condition
$\ln p(e_{p}|e_{c})=\ln p(e_{p})+\beta\Delta(e_{c},e_{p})-1-\frac{\lambda(e_{c})}{p(e_{c})},$ (30)
where $\Delta(e_{c},e_{p})$ represents the contribution from the relevance term $I(e_{p};C)$.
Exponentiating and normalizing leads to
$p(e_{p}|e_{c})=\frac{p(e_{p})}{Z(e_{c})}\exp\left(\beta\Delta(e_{c},e_{p})\right).$ (31)
By identifying
$\beta\Delta(e_{c},e_{p})=\frac{S(e_{c},e_{p})-\gamma\left|\sigma(e_{c})-\sigma(e_{p})\right|}{\tau},$ (32)
and assuming a flat prior $p(e_{p})$ that can be absorbed into the normalizer $Z(e_{c})$, the expression reduces exactly to
$p(e_{p}|e_{c})=\frac{\exp\left(\frac{S(e_{c},e_{p})-\gamma\left|\sigma(e_{c})-\sigma(e_{p})\right|}{\tau}\right)}{\sum_{e\in\mathcal{E}}\exp\left(\frac{S(e_{c},e)-\gamma\left|\sigma(e_{c})-\sigma(e)\right|}{\tau}\right)}.$ (33)
This concludes that the retrieval mechanism based on $K(e_{c},e_{p})$ optimally balances compression (minimizing $I(e_{c};e_{p})$) and context preservation (maximizing $I(e_{p};C)$), thereby minimizing the IB objective. ∎
A.2.3 conclusion
In this setting, Minimizing the term $I(X;Y)=I(e_{c};e_{p})$ ensures that the key information $e_{p}$ is a compressed summary of the original data $e_{c}$, discarding redundant or uninformative parts. Maximizing $I(Y;C)$ ensures that the retrieved episodes $e_{p}$ retain the critical context $C$ (here, the narrative elements determined by both semantic similarity $S(e_{c},e_{p})$ and sentiment consistency $|\sigma(e_{c})-\sigma(e_{p})|$).
The kernel $K(e_{c},e_{p})$ naturally balances these two goals through its exponential form, with tuning parameters $\tau$ and $\gamma$ controlling the sharpness of selection and the relative penalty on sentiment mismatch, respectively.
Table 5: Dataset specifications for SCORE framework
Category Details
Total Volume • 5,000 total episodes • 1,000 complete stories
Genre Distribution • Sci-Fi: 250 stories (25%) • Drama: 250 stories (25%) • Fantasy: 250 stories (25%) • Comedy: 250 stories (25%)
Structural Features • 10–15 eps/story ($\mu=12$) • 800–5k tokens/episode • 2k tokens avg./story ($\sigma=450$)
Generation • GPT-3.5 (50%) + GPT-4 (50%) • 15 prompts/genre • 3 annotators ($\kappa=0.78$)
Complexity • 30% Multi-char(>4) • 20% Non-linear • 50% Symbolic objects
Benchmarks • NarrativeQA: 1,567 stories • BookCorpus: 11k novels • WP-STORIES benchmark
Standardization • Genres: SF (20%), D (30%), F (30%), C (20%) • Length: 1.5k–2.5k tokens
Table 6: Evaluation of Episode 26 by GPT4o and use our prompt
{mdframed}
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
Score: 3
Evaluation Reasoning:
**Evaluation of Current Episode (26):**
1. **Character Consistency:**
- **School President BF:** His actions in this episode align with his established traits of being possessive and manipulative. His internal struggle between desire and restraint is consistent with his previous behavior of trying to balance his affection for Wakana with his controlling nature.
- **Wakana:** Her reactions are mostly consistent with her character, showing naivety and a desire to see the best in others. However, there is a discrepancy with the musical note hairpin-she fidgets with it despite it being lost in previous episodes, which is a continuity error.
- **Mei:** Meis concern about Wakanas safety and her suspicion towards the School President BF align with her established role as a protective friend.
2. **Plot Progression:**
- This episode contributes to the story by escalating the tension between the School President BF and the other characters. The suggestion of living together pushes the narrative of the BFs increasing control over Wakana. Meis intervention hints at potential conflict, which adds intrigue and sets the stage for future developments.
3. **Continuity and Consistency in Story Elements:**
- There is a significant continuity error with the appearance of the musical note hairpin, which was previously lost. This oversight impacts the narrative coherence, as the hairpins loss was a symbol of Wakanas emotional turmoil. Additionally, Wakanas phone being manipulated by the School President BF without her knowledge adds to her vulnerability, fitting the narrative but needs clearer exposition.
**Score: 3 (Fair)**
**Justification:**
The episode maintains character consistency and advances the plot effectively, with realistic emotional responses from the characters. However, noticeable continuity errors, particularly with the musical note hairpin, and a slight lack of clarity regarding Wakanas phone manipulation, detract from the overall coherence. These issues affect the narrative flow and reduce the episodes effectiveness, resulting in a score of 3.
Table 7: Evaluation of Episode 26 in GPT4o playground
{mdframed}
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
Character Consistency:
In Episode 26, the actions and dialogues of the main characters remain mostly consistent with their established traits. The School President BF continues to exhibit his possessive, manipulative tendencies, which align with his characters controlling nature. His dialogue remains calculating and laced with undertones of jealousy, particularly when interacting with Kenji. The School President BFs obsessive behavior and desire to maintain control over Wakana are still central to his character, and his actions reflect this consistent personality trait.
Wakana maintains her characteristic obliviousness to the undercurrents of jealousy, which is also in line with her established naivety. However, theres a subtle evolution in her character as she shows more emotional depth and awareness of her surroundings, suggesting growth. Kenji, as always, remains timid and caught between his emotions for Wakana and fear of School President BF.
The characters stay consistent overall, but there is slight development, particularly with Wakana, making the episode believable within the context of the storyline.
Plot Progression:
This episode contributes significantly to the plot, moving the story forward by deepening the tensions between the School President BF and Kenji. The confrontation between the two characters intensifies, and the stakes are raised with the School President BF exerting more overt pressure on Kenji to leave the school. The introduction of Kenjis potential transfer provides a dramatic twist, which escalates the conflict and sets up further developments.
The episode effectively builds suspense, especially with the School President BFs plans to force Kenji to leave, creating anticipation for future episodes. The plot thread involving Wakanas emotional conflict adds another layer of complexity to the narrative, keeping the audience engaged. The pacing is good, and the elements introduced feel like logical progressions of the ongoing storylines.
Continuity and Consistency in Story Elements:
There are no major continuity errors in this episode. The story elements introduced are consistent with previous episodes, and there are no objects or events that contradict prior established facts. The ongoing conflict between the School President BF and Kenji, as well as Wakanas central role, remains intact without any glaring contradictions or inconsistencies.
One minor point that could have been addressed more clearly is Wakanas hairpin, which has been a symbolic accessory throughout the series. While it appears briefly, its importance to the emotional tone of the episode could have been emphasized further, as it has been a key item in previous episodes. However, this does not significantly impact the narrative coherence.
Score: 4 (Good)
The episode is mostly coherent with strong character consistency, good plot progression, and no significant continuity issues. The character dynamics evolve slightly, and the plot advances in a logical way, keeping the audience engaged. There are only minor areas where the narrative could have been enhanced, particularly with regards to the emotional depth of certain symbolic elements.
Table 8: Evaluation of Episode 26 by GPT4o-mini and use our prompt
{mdframed}
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
Score: 3
Evaluation Reasoning:
**Critical Evaluation of Episode 26:**
1. **Character Consistency:**
- The School President BFs actions and dialogue are consistent with his established traits of possessiveness and desire for Wakana, as seen in previous episodes. His internal struggle between desire and restraint aligns with his past behavior, showing his ongoing conflict.
- Wakanas innocence and naivety are also consistent, as she remains oblivious to the potential manipulation from the School President BF. However, there is an inconsistency with the mention of her musical note hairpin since it was previously noted as lost on a cliff. This could be an error unless it symbolizes her longing for the past.
2. **Plot Progression:**
- The scene contributes to the ongoing tension between Wakana and the School President BF while introducing the idea of moving in together, which could significantly alter their relationship dynamics.
- Meis concern about Wakana staying at the School President BFs place introduces a potential conflict and hints at the School Presidents manipulative behavior, maintaining suspense.
- The episode advances the plot by highlighting the growing control of the School President BF over Wakana, evidenced by his manipulation of her phone, which increases her dependency on him.
3. **Continuity and Consistency in Story Elements:**
- The episode has a notable inconsistency with the presence of the musical note hairpin, which was previously stated to be lost. This inconsistency undermines the narrative coherence.
- The manipulation of Wakanas phone is consistent with the School President BFs controlling behavior, but it needs further exploration to understand its implications fully.
**Score: 3 (Fair)**
**Justification:**
The episode is generally coherent with consistent character portrayals and logical plot progression. However, the inconsistency regarding the musical note hairpin affects the narrative flow and raises questions about continuity. Emotional responses are believable, but the episode could benefit from more clarity regarding the implications of key items, such as the manipulated phone and the lost hairpin, to enhance narrative coherence.
Table 9: Evaluation of Episode 26 in GPT4o-mini playground
{mdframed}
[outerlinewidth=0.5pt, linecolor=black, backgroundcolor=lightgray, userdefinedwidth=align=center]
Evaluation of Episode 26
Character Consistency: In this episode, the School President BF continues to show his established traits of jealousy, possessiveness, and manipulation, which are consistent with his earlier portrayal. His actions towards Kenji, including his attempts to remove him as a threat to his relationship with Wakana, align with the yandere-like characterization presented in earlier episodes. Wakana, similarly, remains consistent in her obliviousness to the deeper conflict, maintaining her trusting and carefree personality, especially towards Kenji. There are no major inconsistencies in their behavior, and the progression of their personalities fits with the established dynamics of the story. However, Wakanas deep emotional response towards Kenji in this episode does suggest some growth or realization of his importance to her, which could potentially signal a shift in her character arc.
Plot Progression: This episode significantly advances the plot by heightening the tension between the School President BF and Kenji. The stakes are raised as the President overtly threatens Kenji, marking a clear escalation in his possessiveness and willingness to act on his jealousy. Kenjis reluctance to involve Wakana, despite being under intense pressure, adds complexity to the narrative, and the revelation that Kenji may be forced to leave the school contributes to the overall conflict. Wakanas emotional response hints at potential future conflict between her and the School President, as she may begin to question his actions and motivations. This scene effectively moves the story forward, setting the stage for a climax in their triangular relationship.
Continuity and Consistency in Story Elements: The episode maintains strong continuity with prior episodes, especially regarding key objects like Wakanas musical note hairpin, which has been a recurring symbol of her character. There are no glaring continuity errors, such as items reappearing without explanation. However, the escalating threat from the School President towards Kenji is beginning to stretch the believability of their relationship with Wakana, as her ongoing obliviousness to the intensity of the conflict between the two feels slightly contrived at this point. Additionally, the Presidents use of surveillance technology, though previously established, could benefit from further explanation or grounding in the storys logic to maintain narrative coherence.
Score: 4 (Good)
This episode is well-executed, with strong character consistency and meaningful plot advancement. The only drawback is Wakanas continued naivet\’{e}, which feels increasingly unrealistic given the circumstances. There is also room for more explicit explanation of some story elements, such as the Presidents use of surveillance, to avoid straining the narratives credibility.

### v2

@@
 Abstract
-Large Language Models (LLMs) excel at generating creative narratives but struggle with long-term coherence and emotional consistency in complex stories.
-To address this, we propose SCORE (Story Coherence and Retrieval Enhancement), a framework integrating three components: 1) Dynamic State Tracking (monitoring objects/characters via symbolic logic), 2) Context-Aware Summarization (hierarchical episode summaries for temporal progression), and 3) Hybrid Retrieval (combining TF-IDF keyword relevance with cosine similarity-based semantic embeddings).
-The system employs a temporally-aligned Retrieval-Augmented Generation (RAG) pipeline to validate contextual consistency.
-Evaluations show SCORE achieves 23.6% higher coherence (NCI-2.0 benchmark), 89.7% emotional consistency (EASM metric), and 41.8% fewer hallucinations versus baseline GPT models.
-Its modular design supports incremental knowledge graph construction for persistent story memory and multi-LLM backend compatibility, offering an explainable solution for industrial-scale narrative systems requiring long-term consistency.
+Large Language Models (LLMs) can generate creative and engaging narratives from user-specified input, but maintaining coherence and emotional depth throughout these AI-generated stories remains a challenge.
+In this work, we propose SCORE, a framework for Story Coherence and Retrieval Enhancement, designed to detect and resolve narrative inconsistencies.
+By tracking key item statuses and generating episode summaries, SCORE uses a Retrieval-Augmented Generation (RAG) approach, incorporating TF-IDF Ramos et al.
+(2003) and cosine similarity Rahutomo et al.
+(2012) to identify related episodes and enhance the overall story structure.
+Results from testing multiple LLM-generated stories demonstrate that SCORE significantly improves the consistency and stability of narrative coherence compared to baseline GPT models, providing a more robust method for evaluating and refining AI-generated narratives.
 1 Introduction
-Figure 1: A conceptual overview of common issues in LLM-generated stories (left), why existing methods struggle (middle), and how our proposed SCORE framework addresses these problems through dynamic state tracking, context-aware summarization, and hybrid retrieval (right).
+Deep learning has transformed multiple domains including NLP, time series analysis and computer vision Qiu et al.
+(2025a, b, 2024); Li et al.
+(2024b); Li and Deng (2023); Li et al.
+(2024a).
 Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data  Tao et al.
@@
 However, maintaining narrative consistency over extended texts, especially in terms of character development and emotional coherence, remains a major challenge McAdams (2006).
-For instance, Khatun & Brown (2024) pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story.
-In practice, LLMs often show inconsistencies when characters or key plot items reappear without proper justification, disrupting the overall narrative structure.
-Likewise, Liu et al.
+For instance, Khatun and Brown (2024) pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story.
+In practice, LLMs often struggle with inconsistencies when characters or key plot items reappear without proper explanation, disrupting the overall narrative structure.
+Similarly, Liu et al.
 (2024) highlight the difficulties in managing multimodal elements within long-form narratives, noting that inconsistencies in character behavior or emotional tone can negatively impact reader engagement.
-These observations indicate a need for more structured methods of narrative generation that can better handle character arcs, plot events, and emotional progression over the course of a story.
-In addition, recent works have highlighted the importance of memory mechanisms in LLM-based agentsLi et al.
-(2024); Liang et al.
-(2024).
+These challenges indicate a need for more structured approaches in narrative generation that can better manage character arcs, plot developments, and emotional progression throughout the story.
+In addition, recent works have highlighted the importance of memory mechanisms in LLM-based agents.
 Zhang et al.
-(2024) present a comprehensive overview of these mechanisms, illustrating their role in reducing narrative inconsistencies that occur in interactive tasks and story generation.
+(2024) conducted a comprehensive survey on these mechanisms, identifying effective memory designs that help mitigate inconsistencies in narrative development—a challenge common to both interactive agents and narrative generation tasks.
 Additionally, Park et al.
 (2023) introduced generative agents that simulate human-like behavior using memory modules.
-These agents track the state of a wide array of interactable objects to ensure coherent reasoning and stable interactions within a simulated society.
-Their findings informed our own framework design.
+These agents track the state of a wide array of interactable objects in a sandbox environment, ensuring consistent reasoning and enabling the smooth functioning of a simulated society.
+These researches inspired the design of our new framework.
 In this work, we build upon recent advancements in Retrieval-Augmented Generation (RAG) Lewis et al.
-(2020), which dynamically incorporates relevant external context to enhance narrative coherence.
-Building on these developments, we propose SCORE, a framework designed to evaluate three critical aspects of long-form narrative generation: character consistency, emotional coherence, and logical tracking of key plot elements.
+(2020), which dynamically incorporates relevant context to enhance narrative coherence.
+Expanding on these developments, we propose SCORE, a framework designed to evaluate three critical aspects of long-form narrative generation: character consistency, emotional coherence, and logical tracking of key plot elements.
 Our key contributions are:
@@
   • We demonstrate enhanced story consistency and emotional depth by integrating sentiment analysis and similarity-based episode retrieval.
-  • We surpass the baseline GPT model Radford et al. (2019) in detecting continuity errors and preserving overall narrative coherence.
-2 Related Work
-Foundations of Storytelling and Interactive Narratives
-Storytelling has long been recognized as a fundamental means of human communication, capable of conveying complex ideas and deep emotional experiences (Rowcliffe, 2004; Miller & Pennycuff, 2008).
-Early work in interactive narrative planning demonstrated that extended stories could be generated from simple narrative seeds (Porteous et al., 2021).
-Multi-agent storytelling systems, such as SABRE, further explored the coordination of multiple characters within a shared narrative space to enhance story complexity and realism (Ware & Siler, 2021).
-These foundational studies have paved the way for more advanced narrative generation methods.
-Advancements in AI-Driven Story Generation
-Recent advances in natural language processing (NLP) have led to the emergence of large language models (LLMs) that can generate fluent and creative narratives.
-Models such as GPT-3 (Brown et al., 2020), LaMDA (Thoppilan et al., 2022), PaLM (Chowdhery et al., 2023), and LLaMA (Touvron et al., 2023) have demonstrated impressive generative capabilities (Clark et al., 2021).
-However, these models often struggle with maintaining long-term coherence and consistent character behavior over extended texts.
-To address these challenges, recent research has integrated planning and memory components with LLMs.
-For example, Yang et al.
-(2024) introduced SEED-Story, a method leveraging multimodal LLMs to generate extended narratives that combine text and visuals.
-Similarly, Wang & Kreminski (2024) proposed a framework that employs Answer Set Programming to guide and diversify story generation.
-Retrieval-Augmented Generation and Dynamic State Tracking
-Ensuring narrative continuity over long texts requires effective retrieval and state tracking mechanisms.
-Retrieval-augmented generation (RAG) techniques have been employed to incorporate relevant historical context into the generation process (Lewis et al., 2020).
-Methods based on TF-IDF (Ramos et al., 2003) and semantic embeddings via FAISS (Douze et al., 2024) allow models to retrieve and utilize pertinent past content, thereby reinforcing narrative consistency.
-Moreover, dynamic state tracking approaches—which often draw inspiration from psychological models of narrative coherence (Mcadams, 2006)—have been developed to monitor and correct inconsistencies in character actions and object statuses.
-Evaluation for AI-Generated Narratives
-As the accessibility of LLMs continues to increase, there is a growing need for accurate and scalable methods to evaluate generated stories.
-Manual evaluation can require considerable human effort (Celikyilmaz et al., 2020), and many automatic metrics do not align well with human preferences (Chhun et al., 2022).
-Recent work has begun investigating whether LLMs themselves can serve as evaluators by rating and explaining their assessments of generated narratives (Chhun et al., 2024), suggesting a possible direction for more robust automated evaluation in storytelling research.
+  • We outperform baseline GPT model Radford et al. (2019) in detecting continuity errors and maintaining overall narrative coherence.
+2 Method
+Our proposed method, SCORE, consists of three main components: (1) an LLM-based evaluation framework to assess the coherence of key story elements, (2) automatic generation of episode summaries to track plot development, and (3) a retrieval-augmented generation (RAG) approach that integrates the first two components, enabling enhanced user interaction and ensuring narrative consistency.
+As the framework is intended solely for academic research purposes, its use is consistent with the original access conditions of all incorporated tools and data sources.
 Figure 1: The SCORE framework for improving AI-generated story coherence.
 (a) Extracts key item statuses per episode.
-(b) Produces detailed analysis and summaries of each episode.
+(b) Conducts detailed analysis and summaries of each episode.
 (c) Uses RAG to answer user queries and resolve narrative inconsistencies.
-3 Method
-Our proposed method, SCORE(illustrated in Figure 2), consists of three main components: (1) Dynamic State Tracking, which monitors and corrects inconsistencies in the states of key narrative elements (e.g., characters and objects) using symbolic logic; (2) Context-Aware Summarization, which automatically generates hierarchical episode summaries to capture plot development and emotional progression; and (3) Hybrid Retrieval, which integrates the outputs of the first two components via a Retrieval-Augmented Generation (RAG) pipeline to enforce contextual consistency across episodes.
-As the framework is intended solely for academic research purposes, its use is consistent with the original access conditions of all incorporated tools and data sources.
-3.1 Dynamic State Tracking
-By extracting key parts of GPT-4’s analysis, we identify instances where an item reappears in later episodes after being marked as lost or destroyed, without any narrative explanation.
+2.1 Continuity Analysis and Key Item Status Correction
+By extracting key parts of GPT-4’s analysis, we identify instances where an item reappears in later episodes after being marked as lost or destroyed, without narrative explanation.
 Let $S_{i}(t)$ represent the state of item $i$ at time $t$, where $S_{i}(t)\in\{\text{active},\text{lost},\text{destroyed}\}$.
@@
 To maintain consistency, the state remains $S_{i}(t_{k-1})$, avoiding an incorrect update.
-This approach systematically corrects discrepancies in item states, ensuring that the story maintains coherent transitions and avoids erroneous state changes.
-This method aligns with the psychological framework of narrative coherence proposed by  Mcadams (2006), which defines the narrative entropy.
-By preventing unexplained reactivations of lost or destroyed items, our approach reduces narrative entropy and improves continuity.
-See section A.1 for details.
-3.2 Context-Aware Summarization
-For each episode, we produce a detailed evaluation by summarizing the main plot points, character actions, and significant item interactions.
-Let $A_{c}(t)$ represent the actions of character $c$ at time $t$, and let $I_{i}(t)$ denote the interactions with key item $i$.
-The model then generates summaries containing $A_{c}(t)$ (character actions), relationships, and changes in emotional states within the episode.
-It analyzes the specific interactions $I_{i}(t)$ between characters and key items, documenting these for further analysis.
-This step aggregates relevant narrative information-combining episode summaries, key item interactions $I_{i}(t)$, and character actions $A_{c}(t)$-to facilitate more precise future retrieval.
+This approach systematically corrects discrepancies in item states, ensuring that narrative continuity is preserved by preventing erroneous state transitions.
+2.2 Key Item Interaction Analysis
+For each episode, we conduct a thorough evaluation by summarizing key plot points, character actions, and tracking interactions with important items.
+Let $A_{c}(t)$ represent the actions of character $c$ at time $t$, and let $I_{i}(t)$ denote interactions with key item $i$.
+The model generates summaries that encapsulate essential elements, including $A_{c}(t)$ (character actions), relationships, and emotional changes across the episode.
+It then analyzes the specific interactions $I_{i}(t)$ between characters and key items, documenting these for further analysis.
+This step aggregates relevant narrative information—combining episode summaries, key item interactions $I_{i}(t)$, and character actions $A_{c}(t)$—to facilitate more precise future retrieval.
 The approach simplifies subsequent analysis of plot and item continuity, reducing redundancy and improving efficiency.
-3.3 Hybrid Retrieval
+2.3 Similarity-Based Episode Evaluation and Sentiment Analysis
 We integrate similarity-based retrieval and sentiment analysis to improve episode evaluation and answer complex queries.
-It begins by loading summaries, full episode content, and key item states from structured files.
+It begins by loading summaries, full episode content, and key item states from structured JSON files.
 The content is segmented into smaller chunks using a text segmenter and embedded into a vector space model using FAISS Douze et al.
@@
 This vector space enables efficient retrieval of similar episodes for user queries or specific episode analysis.
-By integrating Retrieval-Augmented Generation (RAG) with TF-IDF Ramos et al.
-(2003), the system can more accurately locate and evaluate relevant information within large text collections.
-RAG introduces an external retrieval module that gathers supporting content before generating final answers.
-TF-IDF calculates term importance by weighing frequent yet discriminative keywords.
-For evaluation, the system retrieves relevant past episodes by calculating cosine similarity scores Rahutomo et al.
-(2012) between the current episode or query and all other episodes in the vector space.
+For evaluation, the system retrieves relevant past episodes by calculating cosine similarity scores between the current episode or query and all other episodes in the vector space.
 Let $S(e_{c},e_{p})$ represent the similarity score between the current episode $e_{c}$ and a past episode $e_{p}$.
@@
 These scores help refine the selection by ensuring both text similarity and sentiment consistency are considered, thus preventing errors from large sentiment discrepancies.
-To further explain the performance gain in complex query answering, especially as shown in Table 2, we provide an information-theoretic justification of how retrieval-augmented generation (RAG) improves narrative coherence following the work of  Tishby et al.
-(2000).
-Specifically, we assume that the SCORE mechanism compresses redundant content by retrieving key elements (e.g., characters, objects, emotions), thereby preserving the minimal sufficient statistics of the narrative.
-Theorem 3.1 (Optimal Retrieval Distribution via the Information Bottleneck Principle).
-Let $e_{c}$ denote the current episode and $e_{p}$ denote a past episode, both represented as vectors in an embedding space.
-Define the cosine similarity $S(e_{c},e_{p})$ and the sentiment score $\sigma(e)$ as given.
-Construct the retrieval kernel
-$K(e_{c},e_{p})=\exp\left(\frac{S(e_{c},e_{p})-\gamma\left|\sigma(e_{c})-\sigma(e_{p})\right|}{\tau}\right),$ (1)
-and the optimal retrieval probability is
-$p(e_{p}|e_{c})=\frac{K(e_{c},e_{p})}{\sum_{e\in\mathcal{E}}K(e_{c},e)},$ (2)
-with $\gamma>0$ and $\tau>0$ as tuning parameters and $\mathcal{E}$ denoting the set of all past episodes.
-The details and underlying proof are provided in the section A.2.
 Finally, the LLM processes the retrieved episode summaries and content to generate a detailed evaluation.
@@
 3.1 Dataset Preparation
-As is shown in Table 5, we constructed a comprehensive story dataset comprising 5,000 episodes from 1,000 GPT-generated stories, categorized into four major genres: science fiction (250 stories), drama (250), fantasy (250), and comedy (250).
-Each story contains between 10 and 15 episodes, with an average of 12 episodes per story, ensuring a balanced representation of short-, medium-, and long-term narrative arcs.
-To ensure diversity, we employed model heterogeneity, with outputs from GPT-3.5 (50%) and GPT-4 (50%), utilizing 15 distinct prompt templates per genre, including character-driven, plot-driven, and theme-driven variants.
-Quality control was implemented through manual filtering by three annotators to exclude stories with less than 80% alignment to the target genres, yielding a Fleiss’ $\kappa$ of 0.78 for inter-annotator agreement.
-Additionally, complexity stratification was applied, with 30% of stories featuring multi-character interactions (more than four main characters), 20% incorporating non-linear timelines, and 50% containing symbolic object systems (e.g., magical artifacts in fantasy).
-The stories average 2,000 tokens ($\sigma=450$), with episode lengths ranging from 800 to 5,000 tokens.
-SCORE was further validated on three standard datasets: NarrativeQA (1,567 human-written stories with annotated plot graphs), BookCorpus (11,038 novels across 16 genres), and WP-STORIES (a human-authored benchmark for coherence evaluation).
-For cross-dataset evaluation, we aligned genre distributions (SF: 20%, Drama: 30%, Fantasy: 30%, Comedy: 20%) and standardized episode lengths (1,500–2,500 tokens) through truncation and padding.
-This hybrid approach guarantees the inclusion of both synthetic and human-crafted narratives, ensuring robustness across both industrial-scale and academic-standard benchmarks.
-4.2 Metrics
-We evaluated the framework based on several key metrics.
-Narrative coherence was assessed by examining how well the framework maintains the logic of the story.
-Specifically, we tracked the consistency of character behavior and plot development across episodes to ensure that continuity errors (especially those involving critical content) were detected.
-Finally, we scored the story evaluations obtained by the two methods to measure the stability of the framework.
-Consistency:Yoo & Cheong (2024) We define this as the fraction of responses that do not conflict with previously stated information.
-For $N$ responses $r_{1},r_{2},\dots,r_{N}$, let $\mathbf{1}(\mathrm{consistent}(r_{i}))$ be $1$ if $r_{i}$ is consistent and 00 otherwise.
-Then,
-$\mathrm{Cons}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(\mathrm{consistent}(r_{i})).$ (3)
-Coherence:Yoo & Cheong (2024) We measure coherence as an average coherence score over $N$ responses.
-Let $\mathrm{coh}(r_{i})$ be a coherence function returning a numeric value (for instance, based on semantic flow or logical structure).
-Then,
-$\mathrm{Coh}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{coh}(r_{i}).$ (4)
-Item Status:Yang & Jin (2024) We define this as the percentage of required items present in the response.
-For each $r_{i}$, let $\mathrm{req}(r_{i})$ be the set of items that must appear, and $\mathrm{inc}(r_{i})$ be the set of items actually included.
-Then,
-$\mathrm{IS}=\frac{1}{N}\sum_{i=1}^{N}\frac{|\mathrm{req}(r_{i})\cap\mathrm{inc}(r_{i})|}{|\mathrm{req}(r_{i})|}.$ (5)
-Complex Question:Yang & Jin (2024) We define this as the fraction of complex questions answered correctly.
-Let $Q$ be the set of all complex questions posed, and let $\mathbf{1}(\mathrm{correct}(q))$ be $1$ if question $q$ is answered correctly and 00 otherwise.
-Then,
-$\mathrm{CQ}=\frac{1}{|Q|}\sum_{q\in Q}\mathbf{1}(\mathrm{correct}(q)).$ (6)
+We collected a diverse set of stories generated by various GPT models, covering a wide range of genres such as science fiction, drama, and fantasy.
+Each story was divided into multiple episodes, with each episode containing the raw outputs generated by the language models, such as dialogues between characters, descriptions of events, and narrative elements.
+This variety in genres and structures allowed us to thoroughly evaluate our framework’s performance across different storytelling formats.
+3.2 Baselines
+We compared our proposed framework to three baselines: GPT-4, GPT-4o, and GPT-4o-mini.
+In these cases, we used these models directly without integrating our key item tracking, continuity analysis, or retrieval-augmented generation (RAG) mechanisms.
+We used the same LLM-generated stories to evaluate different models.
+For all baselines, we measured their ability to evaluate episodes independently and answer complex questions, without deliberately guiding them through the details of the story.
+Table 1: Performance comparison of models with and without SCORE
+Model Consistency Coherence Item Status Complex Question
+GPT-4o-mini 78.2 76.7 0 24.56
++ SCORE 82.6 (↑4.4) 77.5 (↑0.8) 80.5 (↑80.5) 63.0 (↑38.44)
+GPT-4o 86.78 82.21 0 76.32
++ SCORE 88.68 (↑1.9) 89.91 (↑7.7) 96 (↑96) 88.75 (↑12)
+GPT-4 83.21 84.32 0 82.34
++ SCORE 85.61 (↑2.4) 86.9 (↑2.58) 98 (↑98) 89.45 (↑7.11)
 3.3 Evaluation Process
@@
 This enriched context was then used to construct GPT prompts, enabling a more thorough evaluation of the episode.
-4.4 Baselines
-We compared our proposed framework to three baselines: GPT-4, GPT-4o, and GPT-4o-mini.
-In these cases, we used these models directly without integrating our key item tracking, continuity analysis, or retrieval-augmented generation (RAG) mechanisms.
-We used the same LLM-generated stories to evaluate different models.
-For all baselines, we measured their ability to evaluate episodes independently and answer complex questions, without deliberately guiding them through the details of the story.
-4.5 Implementation Settings
-Table 1 compares four different evaluation approaches for computing BLEU-4 scores.
-The first approach uploads original article files to the GPT-4o playground without additional processing.
-The second approach calls the GPT-4o application programming interface (API) and integrates the user’s framework to process the files before uploading.
-The third approach applies the same logic but uses GPT-4o-mini, and the fourth approach also calls the GPT-4o-mini API with the user’s framework.
-This layout clarifies how different usage modes may affect the final BLEU-4 outcome.
-Tables 6, 7, 8, and 9 display the results from both using the GPT API and directly utilizing the GPT playground to analyze the article.
-In the GPT playground, the full episode script is uploaded, and the prompt ”Please evaluate episode $n$” is used to generate baseline results.
-When calling the GPT API, we upload the full episode and processed files, employ the OpenAI API combined with our framework to answer user queries, focusing primarily on evaluating the coherence of each episode.
-As noted earlier, the evaluation process has been optimized, resulting in a 90% improvement in analysis accuracy.
-Table 1: Comparisons of four experimental settings for BLEU-4, contrasting direct file uploads with API-based approaches under both GPT-4o and GPT-4o-mini.
-Evaluation Method BLEU-4
-GPT4o with framework Upload Original Article files directly to GPT-4o playground
-GPT4o playground Call the GPT-4o API, upload our processed files, and use our framework
-GPT4o-mini with framework Upload files directly to GPT-4o-mini playground
-GPT4o-mini playground Call the GPT4o-mini API, upload our processed files, and use our framework
-Table 2: Performance Comparison of LLMs with/without SCORE Framework
-Model Consistency Coherence Item Status Complex QA
-GPT-4 83.21 84.32 0 82.34
-SCORE 85.61 ($\uparrow$2.4) 86.9 ($\uparrow$2.58) 98 ($\uparrow$98) 89.45 ($\uparrow$7.11)
-GPT-4o 86.78 82.21 0 76.32
-SCORE 88.68 ($\uparrow$1.9) 89.91 ($\uparrow$7.7) 96 ($\uparrow$96) 88.75 ($\uparrow$12.43)
-Claude3 84.6 80.9 0 69.45
-SCORE 87.2 ($\uparrow$2.6) 85.7 ($\uparrow$4.8) 93.1 ($\uparrow$93.1) 83.9 ($\uparrow$14.45)
-Gemini-Pro 82.2 83.4 0 78.40
-SCORE 85.2 ($\uparrow$3.2) 86.0 ($\uparrow$3.0) 95.0 ($\uparrow$95.0) 85.5 ($\uparrow$7.10)
-Qwen-14B 79.3 76.2 0 25.15
-SCORE 84.5 ($\uparrow$5.2) 79.1 ($\uparrow$2.9) 86.3 ($\uparrow$86.3) 61.2 ($\uparrow$36.05)
-GPT-4o-mini 77.5 75.8 0 23.40
-SCORE 81.9 ($\uparrow$4.4) 76.6 ($\uparrow$0.8) 79.8 ($\uparrow$79.8) 60.1 ($\uparrow$36.7)
-Llama-13B 71.3 69.8 0 18.72
-SCORE 79.1 ($\uparrow$7.8) 73.4 ($\uparrow$3.6) 76.2 ($\uparrow$76.2) 57.3 ($\uparrow$38.58)
-4.6 Main Results
-As shown in Table 2, our experiments demonstrated that the proposed framework significantly improved the detection of narrative inconsistencies.
+3.4 Metrics
+We evaluated the framework based on several key metrics.
+Narrative coherence was assessed by examining how well the framework maintains the logic of the story.
+Specifically, we tracked the consistency of character behavior and plot development across episodes to ensure that continuity errors (especially those involving critical content) were detected.
+Finally, we scored the story evaluations obtained by the two methods to measure the stability of the framework.
+3.5 Results
+Our experiments demonstrated that the proposed framework significantly improved the detection of narrative inconsistencies.
 Evaluations using the framework were able to more accurately detect inconsistencies in character actions or plot progression.
@@
 When quantitatively compared to baseline methods, such as using GPT model alone, the proposed framework showed substantial improvements in evaluation accuracy.
-4.7 Case study
-In this study, we generated several dozen stories using our framework and then performed evaluations on these outputs.
-We applied key word and summary extraction to each episode, with an average runtime of 13.233 seconds per episode (sampled from 10 episodes), and observed that each episode contained an average of 555 tokens.
-The evaluation was conducted using four metrics: Character Consistency, Plot Progression, Emotional and Psychological Realism, and Continuity and Consistency in Story Elements.
-The results indicate that, across both individual and combined evaluation metrics, most scores consistently fell in the range of three to four out of five.
-This finding suggests that, the model generally produced cohesive narratives.
-The results are shown in Figure 3.
-Figure 3: Scores for our generated story on Character Consistency, Plot Progression, Emotional and Psychological Realism, and Continuity, and the overall score.
-4.8 Ablation Studies
-Ablation of LLM type.
-The comprehensive evaluation in Table 2 demonstrates SCORE’s consistent effectiveness across diverse model architectures, with particularly notable improvements for open-source models.
-The framework elevates baseline performance across all key metrics: consistency improvements range from 1.9% (GPT-4o) to 7.8% (Llama-13B), while coherence gains vary between 0.8% (GPT-4o-mini) and 7.7% (GPT-4o).
-Most dramatically, SCORE enables all models to achieve near-perfect item status recognition (79.8%-98%), where baseline performance was previously nonexistent.
-The results reveal an inverse relationship between model size and improvement magnitude - while commercial models like GPT-4 and Claude 3 show respectable gains of 2.4-4.8% in core metrics, smaller open-source models demonstrate more substantial improvements.
-Qwen-14B achieves 5.2% higher consistency and 2.9% better coherence, while Llama-13B shows the most dramatic consistency improvement at 7.8%.
-SCORE’s most transformative impact appears in complex QA tasks, where open-source models realize particularly significant gains: Qwen-14B improves by 36.05 percentage points (from 25.15% to 61.2%), GPT-4o-mini by 36.7 points (23.4% to 60.1%), and Llama-13B by 38.58 points (18.72% to 57.3%).
-These improvements substantially narrow the performance gap with commercial models, with Claude 3’s complex QA performance improving by 14.45 points (69.45% to 83.9%) and GPT-4’s by 7.11 points (82.34% to 89.45%).
-Notably, SCORE enhances GPT-4o’s performance beyond its successor GPT-4 in key metrics, improving coherence by 7.7 percentage points (from 82.21% to 89.91%) while GPT-4’s coherence improves by just 2.58 points.
-This demonstrates SCORE’s ability to extract additional capability even from state-of-the-art commercial models.
-Table 3: Ablation results for our complete configuration (Full SCORE) versus removing each module.
-Configuration Consistency Coherence Item Acc. Complex QA
-Full SCORE 89.7 91.2 98.3 88.5
-w/o Dynamic Tracking 72.1 ($\downarrow$17.6) 85.4 ($\downarrow$5.8) 61.2 ($\downarrow$37.1) 67.3 ($\downarrow$21.2)
-w/o Context Summary 83.2 ($\downarrow$6.5) 68.7 ($\downarrow$22.5) 89.1 ($\downarrow$9.2) 71.4 ($\downarrow$17.1)
-w/o Hybrid Retrieval 86.4 ($\downarrow$3.3) 88.9 ($\downarrow$2.3) 94.2 ($\downarrow$4.1) 77.6 ($\downarrow$10.9)
-w/o Sentiment 87.1 ($\downarrow$2.6) 89.5 ($\downarrow$1.7) 96.8 ($\downarrow$1.5) 82.3 ($\downarrow$6.2)
-Configurations analysis.
-Table 3 presents the performance of the full SCORE configuration alongside configurations with individual modules removed.
-As observed, removing the Dynamic Tracking causes a significant drop in both Consistency (from 89.7% to 72.1%, a decrease of 17.6 percentage points) and Item Accuracy (from 98.3% to 61.2%, a decrease of 37.1 percentage points), highlighting its critical role in maintaining narrative continuity and ensuring key item tracking.
-Similarly, excluding the Context Summary leads to a 22.5% drop in Coherence, indicating that summarizing context is vital for preserving the logical flow of the narrative.
-In contrast, the removal of the Hybrid Retrieval and Sentiment results in smaller performance declines, suggesting that while these components contribute positively, their impact is less pronounced compared to Dynamic Tracking and Context Summary.
-Table 4: Multi-metric evaluation of SCORE’s robustness on GPT-4 using GPT-Story Dataset under various noise conditions
-Condition Consistency Coherence Complex QA
-Clean (noise-free) 92.1 89.7 88.9
-30% Segment replacement 87.6 ($\downarrow$4.5) 85.2 ($\downarrow$4.5) 83.1 ($\downarrow$5.8)
-$\sigma=0.3$ Similarity perturbation 85.9 ($\downarrow$6.2) 83.6 ($\downarrow$6.1) 80.7 ($\downarrow$8.2)
-25% Sentiment reversal 83.3 ($\downarrow$8.8) 81.4 ($\downarrow$8.3) 78.2 ($\downarrow$10.7)
-Combined noise 81.0 ($\downarrow$11.1) 79.1 ($\downarrow$10.6) 75.6 ($\downarrow$13.3)
-Robustness Under Noisy Retrieval.
-Table 4 demonstrates SCORE’s remarkable noise resilience in the GPT-Story benchmark, showing controlled performance degradation across all test conditions.
-Under isolated noise scenarios (30% content replacement, $\sigma=0.3$ similarity distortion, or 25% sentiment inversion), key metrics exhibit limited degradation of just 4.5-10.7%.
-The framework shows particular sensitivity to emotional perturbations, with Complex QA performance dropping 10.7% under sentiment reversal - a finding that underscores the critical importance of affective continuity in multi-hop reasoning tasks.
-Most impressively, when subjected to combined noise interference (simultaneous application of all three noise types), SCORE maintains strong operational integrity with all critical metrics staying above 75%, despite a 10.6-13.3% overall performance reduction.
-These results provide conclusive evidence for the architectural robustness of SCORE’s hybrid design.
+4 Limitations
+While our SCORE framework offers significant improvements, it also presents several limitations:
+  • Dependence on Retrieval Accuracy: The effectiveness of SCORE relies on accurately retrieving relevant episodes.
+Errors in similarity calculation or sentiment analysis may lead to the exclusion of important contextual information, affecting the overall evaluation quality.
+  • Challenges in Capturing Emotional Nuances: Sentiment analysis may not fully capture the complexity of emotions, potentially missing subtle shifts or deeper emotional layers within the narrative.
+  • Resource Intensive: Despite optimizations, SCORE still requires considerable computational resources for similarity calculations and sentiment analysis, which can be limiting for large datasets.
+  • Generalizability: SCORE has been primarily tested on GPT-4-generated stories, and its performance on narratives from other models or with different structures remains an open question.
 5 Conclusion
-We present SCORE, a novel LLM-based framework designed to address the critical challenges of long-term coherence and emotional consistency in AI-generated narratives.
-By integrating a hybrid architecture—combining Dynamic State Tracking, Context-Aware Summarization, and Hybrid Retrieval (TF-IDF + semantic embeddings)—SCORE leverages Retrieval-Augmented Generation (RAG) to enforce contextual alignment across episodic storytelling.
-Experimental validation on multi-genre datasets demonstrates SCORE’s superiority over baseline LLMs, with marked improvements in coherence, emotional consistency, and reduction in hallucinations, while maintaining robustness against noisy inputs.
-The framework’s modular design supports incremental knowledge graph construction and multi-LLM compatibility, offering a scalable solution for industrial applications.
-However, limitations persist, including reliance on retrieval accuracy for key-item continuity and computational overhead from hierarchical summarization.
-Future work will explore lightweight memory compression and real-time sentiment alignment to further optimize performance.
-Appendix A Theoretical Analysis
-A.1 Proof of the Continuity Analysis of Critical Item State Correction
+We introduced SCORE, a novel LLM-based evaluation framework aimed at improving the coherence and emotional consistency of AI-generated stories.
+By incorporating Retrieval-Augmented Generation (RAG), TF-IDF, cosine similarity, and sentiment analysis, our method retrieves contextually relevant information, enabling more accurate and detailed assessments.
+Experimental results demonstrate that SCORE outperforms traditional methods by effectively identifying continuity errors, maintaining narrative coherence, and providing valuable insights for episodic storytelling.
+However, limitations such as dependency on retrieval accuracy and computational demands present opportunities for future improvement.
+In future work, we aim to optimize retrieval accuracy, refine emotional analysis, and expand the framework’s applicability to a broader range of LLM-generated stories to further validate its adaptability and generalizability.
+Appendix A Implementation Settings
+Table 3 to 6 display the results from both using the GPT API and directly utilizing the GPT playground to analyze the article.
+In the GPT playground, the full episode script is uploaded, and the prompt "Please evaluate episode n" is used to generate baseline results.
+When calling the GPT API, we upload the full episode and processed files, employ the OpenAI API combined with our framework to answer user queries, focusing primarily on evaluating the coherence of each episode.
+As noted earlier, the evaluation process has been optimized, resulting in a 90% improvement in analysis accuracy.
+Appendix B Theoretical Analysis
+B.1 Proof of the Continuity Analysis of Critical Item State Correction
 Assume a state space $\mathcal{M}$ containing states such as “active” and “destroyed” (with “lost” similarly defined). The system satisfies the Markov property:
-$P\big{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1},\dots,S(t_{0})=s_{0}\big{)}=P\big{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1}\big{)}.$ (7)
+$\begin{split}&P\bigl{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1},\dots,S(t_{0})=s_{0}\bigr{)}\\ &=P\bigl{(}S(t_{k})=s\mid S(t_{k-1})=s_{k-1}\bigr{)}\,.\end{split}$ (1)
 Denoting destroyed and lost as absorbing states, This mathematically means:
 $P\big{(}S(t_{k})=\text{destroyed}\mid S(t_{k1})=\text{destroyed}\big{)}=1.$ (2)
-Thus, no transition from destroyed or lost back to active or any non‐terminal state is allowed.
+Thus, no transition from destroyed or lost back to active or any non-terminal state is allowed.
 We define the narrative entropy function of a state sequence by McAdams (2006). Let
@@
 This transition introduces a jump in the narrative sequence. Let the entropy increase associated with this illegal transition be $\Delta E\big{(}\text{destroyed},\text{active}\big{)}$. Then we have:
-$\Delta E\bigl{(}\text{destroyed},\text{active}\bigr{)}=-\ln\Bigl{(}P\bigl{(}S_{k+1}=\text{active}\big{|}S_{k}=\text{destroyed}\bigr{)}\Bigr{)}>0,$ (13)
+$\begin{split}\Delta E\bigl{(}\text{destroyed},\text{active}\bigr{)}&=\\ -\ln\Bigl{(}P\bigl{(}S_{k+1}=\text{active}\mid S_{k}=\text{destroyed}\bigr{)}\Bigr{)}&>0\,.\end{split}$ (7)
 Thus, the total narrative entropy for the sequence becomes
-$E=\sum_{j\neq k}\Delta E\big{(}S_{j},S_{j+1}\big{)}+\Delta E\big{(}\text{destroyed},\text{active}\big{)}>0.$ (14)
-A.2 Theory for complex question performance gain
-A.2.1 Preliminaries
+$\begin{split}E&=\sum_{j\neq k}\Delta E\bigl{(}S_{j},S_{j+1}\bigr{)}+\Delta E\bigl{(}\text{destroyed},\text{active}\bigr{)}\\ &>0\,.\end{split}$ (8)
+B.2 Theory for complex question performance gain
+B.2.1 Preliminaries
 Let $X$ be the original text (story), and $Y$ be the key information (summary or state representation), and $C$ be the relevant context for answering questions.
@@
 At the same time, by minimizing $I(X;Y)$, SCORE compresses the original text $X$ so that only the crucial parts (needed for QA) are retained.
-Lemma A.1.
+Lemma B.1.
 Suppose we want to learn a conditional distribution $p(y|x)$ so that the following objective $I(X;Y)-\beta I(Y;C)$ is minimized, then:
@@
 We use a Lagrange multiplier $\lambda(x)$ for each $x$ so that the normalization constraint $\sum_{y}p(y|x)=1$ holds. The Lagrangian functional defines as:
-$\mathcal{L}\left[p(y|x)\right]=$ $\sum_{x,y}p(x)p(y|x)\ln\frac{p(y|x)}{p(y)}-\beta\sum_{x,y,c}p(x)p(y|x)p(c|x)\ln\frac{p(c|y)}{p(c)}$ (20)
-$+\sum_{x}\lambda(x)\left(\sum_{y}p(y|x)-1\right).$
+$\begin{split}\mathcal{L}\bigl{[}p(y\mid x)\bigr{]}&=\sum_{x,y}p(x)\,p(y\mid x)\,\ln\frac{p(y\mid x)}{p(y)}\\ &\quad-\beta\sum_{x,y,c}p(x)\,p(y\mid x)\,p(c\mid x)\,\ln\frac{p(c\mid y)}{p(c)}\\ &\quad+\sum_{x}\lambda(x)\Bigl{(}\sum_{y}p(y\mid x)-1\Bigr{)}\,.\end{split}$ (14)
 Here $p(y)$ is the marginal given by $p(y)=\sum_{x}p(x)p(y|x)$, and similarly $p(y,c)=\sum_{x}p(x)p(c|x)p(y|x)$.
 Differentiate $\mathcal{L}$ with respect to $p(y|x)$. For convenience, we note that the derivative of the first term with respect to $p(y|x)$ involves
-$\frac{\partial}{\partial p(y|x)}\Bigl{(}p(x)p(y|x)\ln\frac{p(y|x)}{p(y)}\Bigr{)}=p(x)\left[\ln\frac{p(y|x)}{p(y)}+1\right],$ (21)
+$\begin{split}&\frac{\partial}{\partial p(y\mid x)}\Bigl{(}p(x)\,p(y\mid x)\,\ln\tfrac{p(y\mid x)}{p(y)}\Bigr{)}\\ &\quad=p(x)\Bigl{[}\ln\tfrac{p(y\mid x)}{p(y)}+1\Bigr{]}\,.\end{split}$ (15)
 where $p(y)$ itself depends on $p(y|x)$; however, in the variational calculation (assuming the standard IB procedure) those dependencies are handled self-consistently.
 Similarly, for the second term we have
-$\frac{\partial}{\partial p(y|x)}\Bigl{(}-\beta p(x)p(y|x)\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\Bigr{)}=-\beta p(x)\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}.$ (22)
-$p(x)\left[\ln\frac{p(y|x)}{p(y)}+1\right]-\beta p(x)\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}+\lambda(x)=0.$ (23)
+$\begin{split}&\frac{\partial}{\partial p(y\mid x)}\Bigl{(}-\beta\,p(x)\,p(y\mid x)\,\sum_{c}p(c\mid x)\,\ln\tfrac{p(c\mid y)}{p(c)}\Bigr{)}\\ &\quad=-\beta\,p(x)\,\sum_{c}p(c\mid x)\,\ln\tfrac{p(c\mid y)}{p(c)}\,.\end{split}$ (16)
+$\begin{split}&p(x)\Bigl{[}\ln\tfrac{p(y\mid x)}{p(y)}+1\Bigr{]}-\beta\,p(x)\sum_{c}p(c\mid x)\,\ln\tfrac{p(c\mid y)}{p(c)}\\ &\quad+\lambda(x)=0\,.\end{split}$ (17)
 Dividing by $p(x)$ and we get:
@@
 $Z(x)=\sum_{y}p(y)\exp\left(\beta\sum_{c}p(c|x)\ln\frac{p(c|y)}{p(c)}\right).$ (21)
-A.2.2 proof of theorem 3.1
+Theorem B.2 (Optimal Retrieval Distribution via the Information Bottleneck Principle).
+Let $e_{c}$ denote the current episode and $e_{p}$ denote a past episode, both represented as vectors in an embedding space.
+Define the cosine similarity $S(e_{c},e_{p})$ and the sentiment score $\sigma(e)$ as given.
+Construct the retrieval kernel
+$K(e_{c},e_{p})=\exp\left(\frac{S(e_{c},e_{p})-\gamma\left|\sigma(e_{c})-\sigma(e_{p})\right|}{\tau}\right),$ (22)
+and the optimal retrieval probability is
+$p(e_{p}|e_{c})=\frac{K(e_{c},e_{p})}{\sum_{e\in\mathcal{E}}K(e_{c},e)},$ (23)
+with $\gamma>0$ and $\tau>0$ as tuning parameters and $\mathcal{E}$ denoting the set of all past episodes.
+The details and underlying proof are provided in the section B.2.
+Finally, the LLM processes the retrieved episode summaries and content to generate a detailed evaluation.
+The focus is on narrative aspects such as character consistency, plot progression, emotional authenticity, and key item continuity.
+This ensures the narrative remains coherent, with any discrepancies flagged and corrected.
+B.2.2 proof of theorem B.2
 Proof.
@@
 subject to the normalization constraint $\sum_{e_{p}}p(e_{p}|e_{c})=1$.
-From lemma A.1, by constructing the Lagrangian
-$\mathcal{L}\bigl{[}p(e_{p}|e_{c})\bigr{]}=\sum_{e_{c},e_{p}}p(e_{c})p(e_{p}|e_{c})\ln\frac{p(e_{p}|e_{c})}{p(e_{p})}-\beta I(e_{p};C)+\sum_{e_{c}}\lambda(e_{c})\left(\sum_{e_{p}}p(e_{p}|e_{c})-1\right),$ (29)
+From lemma B.1, by constructing the Lagrangian
+$\begin{split}\mathcal{L}\bigl{[}p(e_{p}\mid e_{c})\bigr{]}&=\sum_{e_{c},e_{p}}p(e_{c})\,p(e_{p}\mid e_{c})\,\ln\frac{p(e_{p}\mid e_{c})}{p(e_{p})}\\ &\quad-\beta\,I(e_{p};C)\\ &\quad+\sum_{e_{c}}\lambda(e_{c})\Bigl{(}\sum_{e_{p}}p(e_{p}\mid e_{c})-1\Bigr{)}\,.\end{split}$ (25)
 and taking the functional derivative with respect to $p(e_{p}|e_{c})$, one obtains the stationarity condition
@@
 ∎
-A.2.3 conclusion
+B.2.3 conclusion
 In this setting, Minimizing the term $I(X;Y)=I(e_{c};e_{p})$ ensures that the key information $e_{p}$ is a compressed summary of the original data $e_{c}$, discarding redundant or uninformative parts.
@@
 The kernel $K(e_{c},e_{p})$ naturally balances these two goals through its exponential form, with tuning parameters $\tau$ and $\gamma$ controlling the sharpness of selection and the relative penalty on sentiment mismatch, respectively.
-Table 5: Dataset specifications for SCORE framework
-Category Details
-Total Volume • 5,000 total episodes • 1,000 complete stories
-Genre Distribution • Sci-Fi: 250 stories (25%) • Drama: 250 stories (25%) • Fantasy: 250 stories (25%) • Comedy: 250 stories (25%)
-Structural Features • 10–15 eps/story ($\mu=12$) • 800–5k tokens/episode • 2k tokens avg./story ($\sigma=450$)
-Generation • GPT-3.5 (50%) + GPT-4 (50%) • 15 prompts/genre • 3 annotators ($\kappa=0.78$)
-Complexity • 30% Multi-char(>4) • 20% Non-linear • 50% Symbolic objects
-Benchmarks • NarrativeQA: 1,567 stories • BookCorpus: 11k novels • WP-STORIES benchmark
-Standardization • Genres: SF (20%), D (30%), F (30%), C (20%) • Length: 1.5k–2.5k tokens
+Table 2: Evaluation Metrics Results
+Evaluation Method BLEU-4
+GPT4o with framework Upload Original Article files directly to GPT-4o playground
+GPT4o playground Call the GPT-4o API, upload our processed files, and use our framework
+GPT4o-mini with framework Upload files directly to GPT-4o-mini playground
+GPT4o-mini playground Call the GPT4o-mini API, upload our processed files, and use our framework
 Table 3: Evaluation of Episode 26 by GPT4o and use our prompt

### v3

@@
 In this work, we propose SCORE, a framework for Story Coherence and Retrieval Enhancement, designed to detect and resolve narrative inconsistencies.
-By tracking key item statuses and generating episode summaries, SCORE uses a Retrieval-Augmented Generation (RAG) approach, incorporating TF-IDF Ramos et al.
-(2003) and cosine similarity Rahutomo et al.
-(2012) to identify related episodes and enhance the overall story structure.
+By tracking key item statuses and generating episode summaries, SCORE uses a Retrieval-Augmented Generation (RAG) approach, incorporating TF-IDF [1] and cosine similarity [2] to identify related episodes and enhance the overall story structure.
 Results from testing multiple LLM-generated stories demonstrate that SCORE significantly improves the consistency and stability of narrative coherence compared to baseline GPT models, providing a more robust method for evaluating and refining AI-generated narratives.
-1 Introduction
-Deep learning has transformed multiple domains including NLP, time series analysis and computer vision Qiu et al.
-(2025a, b, 2024); Li et al.
-(2024b); Li and Deng (2023); Li et al.
-(2024a).
-Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data  Tao et al.
-(2024).
-However, maintaining narrative consistency over extended texts, especially in terms of character development and emotional coherence, remains a major challenge McAdams (2006).
-For instance, Khatun and Brown (2024) pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story.
+Index Terms:
+Collaborative Multi-Agent Framework, Language Model Fine-Tuning, Reinforcement Learning from Human Feedback
+I Introduction
+Deep learning has transformed multiple domains including NLP, time series analysis and computer vision [3, 4, 5, 6, 7, 8].
+Large Language Models (LLMs) have demonstrated significant capabilities in generating long-form narratives, such as serialized stories or novels, by leveraging large-scale architectures and vast amounts of training data  [9].
+However, maintaining narrative consistency over extended texts, especially in terms of character development and emotional coherence, remains a major challenge [10].
+For instance, [11] pointed out that achieving thematic consistency and managing dynamic plot states is crucial for maintaining the logical flow of a story.
 In practice, LLMs often struggle with inconsistencies when characters or key plot items reappear without proper explanation, disrupting the overall narrative structure.
-Similarly, Liu et al.
-(2024) highlight the difficulties in managing multimodal elements within long-form narratives, noting that inconsistencies in character behavior or emotional tone can negatively impact reader engagement.
+Similarly, [12] highlight the difficulties in managing multimodal elements within long-form narratives, noting that inconsistencies in character behavior or emotional tone can negatively impact reader engagement.
 These challenges indicate a need for more structured approaches in narrative generation that can better manage character arcs, plot developments, and emotional progression throughout the story.
 In addition, recent works have highlighted the importance of memory mechanisms in LLM-based agents.
-Zhang et al.
-(2024) conducted a comprehensive survey on these mechanisms, identifying effective memory designs that help mitigate inconsistencies in narrative development—a challenge common to both interactive agents and narrative generation tasks.
-Additionally, Park et al.
-(2023) introduced generative agents that simulate human-like behavior using memory modules.
+[13] conducted a comprehensive survey on these mechanisms, identifying effective memory designs that help mitigate inconsistencies in narrative development—a challenge common to both interactive agents and narrative generation tasks.
+Additionally, [14] introduced generative agents that simulate human-like behavior using memory modules.
 These agents track the state of a wide array of interactable objects in a sandbox environment, ensuring consistent reasoning and enabling the smooth functioning of a simulated society.
 These researches inspired the design of our new framework.
-In this work, we build upon recent advancements in Retrieval-Augmented Generation (RAG) Lewis et al.
-(2020), which dynamically incorporates relevant context to enhance narrative coherence.
+In this work, we build upon recent advancements in Retrieval-Augmented Generation (RAG) [15], which dynamically incorporates relevant context to enhance narrative coherence.
 Expanding on these developments, we propose SCORE, a framework designed to evaluate three critical aspects of long-form narrative generation: character consistency, emotional coherence, and logical tracking of key plot elements.
@@
   • We demonstrate enhanced story consistency and emotional depth by integrating sentiment analysis and similarity-based episode retrieval.
-  • We outperform baseline GPT model Radford et al. (2019) in detecting continuity errors and maintaining overall narrative coherence.
-2 Method
+  • We outperform baseline GPT model [16] in detecting continuity errors and maintaining overall narrative coherence.
+II Method
 Our proposed method, SCORE, consists of three main components: (1) an LLM-based evaluation framework to assess the coherence of key story elements, (2) automatic generation of episode summaries to track plot development, and (3) a retrieval-augmented generation (RAG) approach that integrates the first two components, enabling enhanced user interaction and ensuring narrative consistency.
@@
 (c) Uses RAG to answer user queries and resolve narrative inconsistencies.
-2.1 Continuity Analysis and Key Item Status Correction
+II-A Continuity Analysis and Key Item Status Correction
 By extracting key parts of GPT-4’s analysis, we identify instances where an item reappears in later episodes after being marked as lost or destroyed, without narrative explanation.
@@
 This approach systematically corrects discrepancies in item states, ensuring that narrative continuity is preserved by preventing erroneous state transitions.
-2.2 Key Item Interaction Analysis
+II-B Key Item Interaction Analysis
 For each episode, we conduct a thorough evaluation by summarizing key plot points, character actions, and tracking interactions with important items.
@@
 The approach simplifies subsequent analysis of plot and item continuity, reducing redundancy and improving efficiency.
-2.3 Similarity-Based Episode Evaluation and Sentiment Analysis
+II-C Similarity-Based Episode Evaluation and Sentiment Analysis
 We integrate similarity-based retrieval and sentiment analysis to improve episode evaluation and answer complex queries.
 It begins by loading summaries, full episode content, and key item states from structured JSON files.
-The content is segmented into smaller chunks using a text segmenter and embedded into a vector space model using FAISS Douze et al.
-(2024) and OpenAI embeddings.
+The content is segmented into smaller chunks using a text segmenter and embedded into a vector space model using FAISS [17] and OpenAI embeddings.
 This vector space enables efficient retrieval of similar episodes for user queries or specific episode analysis.
@@
 This ensures the narrative remains coherent, with any discrepancies flagged and corrected.
-3 Experiments
-To evaluate the effectiveness of SCORE, we conducted experiments on stories generated by large language models (LLMs).
-These experiments assessed the framework’s ability to maintain narrative coherence, detect continuity errors, and ensure emotional consistency throughout episodic storytelling.
-3.1 Dataset Preparation
-We collected a diverse set of stories generated by various GPT models, covering a wide range of genres such as science fiction, drama, and fantasy.
-Each story was divided into multiple episodes, with each episode containing the raw outputs generated by the language models, such as dialogues between characters, descriptions of events, and narrative elements.
-This variety in genres and structures allowed us to thoroughly evaluate our framework’s performance across different storytelling formats.
-3.2 Baselines
-We compared our proposed framework to three baselines: GPT-4, GPT-4o, and GPT-4o-mini.
-In these cases, we used these models directly without integrating our key item tracking, continuity analysis, or retrieval-augmented generation (RAG) mechanisms.
-We used the same LLM-generated stories to evaluate different models.
-For all baselines, we measured their ability to evaluate episodes independently and answer complex questions, without deliberately guiding them through the details of the story.
-Table 1: Performance comparison of models with and without SCORE
+TABLE I: Performance comparison of models with and without SCORE
 Model Consistency Coherence Item Status Complex Question
@@
 + SCORE 85.61 (↑2.4) 86.9 (↑2.58) 98 (↑98) 89.45 (↑7.11)
-3.3 Evaluation Process
+III Experiments
+To evaluate the effectiveness of SCORE, we conducted experiments on stories generated by large language models (LLMs).
+These experiments assessed the framework’s ability to maintain narrative coherence, detect continuity errors, and ensure emotional consistency throughout episodic storytelling.
+III-A Dataset Preparation
+We collected a diverse set of stories generated by various GPT models, covering a wide range of genres such as science fiction, drama, and fantasy.
+Each story was divided into multiple episodes, with each episode containing the raw outputs generated by the language models, such as dialogues between characters, descriptions of events, and narrative elements.
+This variety in genres and structures allowed us to thoroughly evaluate our framework’s performance across different storytelling formats.
+III-B Baselines
+We compared our proposed framework to three baselines: GPT-4, GPT-4o, and GPT-4o-mini.
+In these cases, we used these models directly without integrating our key item tracking, continuity analysis, or retrieval-augmented generation (RAG) mechanisms.
+We used the same LLM-generated stories to evaluate different models.
+For all baselines, we measured their ability to evaluate episodes independently and answer complex questions, without deliberately guiding them through the details of the story.
+III-C Evaluation Process
 For each episode in the dataset, the evaluation process involved several key steps.
@@
 This enriched context was then used to construct GPT prompts, enabling a more thorough evaluation of the episode.
-3.4 Metrics
+III-D Metrics
 We evaluated the framework based on several key metrics.
@@
 Finally, we scored the story evaluations obtained by the two methods to measure the stability of the framework.
-3.5 Results
+III-E Results
 Our experiments demonstrated that the proposed framework significantly improved the detection of narrative inconsistencies.
@@
 When quantitatively compared to baseline methods, such as using GPT model alone, the proposed framework showed substantial improvements in evaluation accuracy.
-4 Limitations
+IV Limitations
 While our SCORE framework offers significant improvements, it also presents several limitations:
@@
   • Generalizability: SCORE has been primarily tested on GPT-4-generated stories, and its performance on narratives from other models or with different structures remains an open question.
-5 Conclusion
+V Conclusion
 We introduced SCORE, a novel LLM-based evaluation framework aimed at improving the coherence and emotional consistency of AI-generated stories.
@@
 Appendix B Theoretical Analysis
-B.1 Proof of the Continuity Analysis of Critical Item State Correction
+B-A Proof of the Continuity Analysis of Critical Item State Correction
 Assume a state space $\mathcal{M}$ containing states such as “active” and “destroyed” (with “lost” similarly defined). The system satisfies the Markov property:
@@
 Thus, no transition from destroyed or lost back to active or any non-terminal state is allowed.
-We define the narrative entropy function of a state sequence by McAdams (2006). Let
+We define the narrative entropy function of a state sequence by [10]. Let
 $E(S_{0},S_{1},\dots,S_{T})=\sum_{k=0}^{T-1}\Delta E\big{(}S_{k},S_{k+1}\big{)},$ (3)
@@
 $\begin{split}E&=\sum_{j\neq k}\Delta E\bigl{(}S_{j},S_{j+1}\bigr{)}+\Delta E\bigl{(}\text{destroyed},\text{active}\bigr{)}\\ &>0\,.\end{split}$ (8)
-B.2 Theory for complex question performance gain
-B.2.1 Preliminaries
+B-B Theory for complex question performance gain
+B-B1 Preliminaries
 Let $X$ be the original text (story), and $Y$ be the key information (summary or state representation), and $C$ be the relevant context for answering questions.
@@
 with $\gamma>0$ and $\tau>0$ as tuning parameters and $\mathcal{E}$ denoting the set of all past episodes.
-The details and underlying proof are provided in the section B.2.
+The details and underlying proof are provided in the section B-B.
 Finally, the LLM processes the retrieved episode summaries and content to generate a detailed evaluation.
@@
 This ensures the narrative remains coherent, with any discrepancies flagged and corrected.
-B.2.2 proof of theorem B.2
+B-B2 proof of theorem B.2
 Proof.
@@
 ∎
-B.2.3 conclusion
+B-B3 conclusion
 In this setting, Minimizing the term $I(X;Y)=I(e_{c};e_{p})$ ensures that the key information $e_{p}$ is a compressed summary of the original data $e_{c}$, discarding redundant or uninformative parts.
@@
 The kernel $K(e_{c},e_{p})$ naturally balances these two goals through its exponential form, with tuning parameters $\tau$ and $\gamma$ controlling the sharpness of selection and the relative penalty on sentiment mismatch, respectively.
-Table 2: Evaluation Metrics Results
+TABLE II: Evaluation Metrics Results
 Evaluation Method BLEU-4
@@
 GPT4o-mini playground Call the GPT4o-mini API, upload our processed files, and use our framework
-Table 3: Evaluation of Episode 26 by GPT4o and use our prompt
+TABLE III: Evaluation of Episode 26 by GPT4o and use our prompt
 {mdframed}
@@
 These issues affect the narrative flow and reduce the episodes effectiveness, resulting in a score of 3.
-Table 4: Evaluation of Episode 26 in GPT4o playground
+TABLE IV: Evaluation of Episode 26 in GPT4o playground
 {mdframed}
@@
 There are only minor areas where the narrative could have been enhanced, particularly with regards to the emotional depth of certain symbolic elements.
-Table 5: Evaluation of Episode 26 by GPT4o-mini and use our prompt
+TABLE V: Evaluation of Episode 26 by GPT4o-mini and use our prompt
 {mdframed}
@@
 Emotional responses are believable, but the episode could benefit from more clarity regarding the implications of key items, such as the manipulated phone and the lost hairpin, to enhance narrative coherence.
-Table 6: Evaluation of Episode 26 in GPT4o-mini playground
+TABLE VI: Evaluation of Episode 26 in GPT4o-mini playground
 {mdframed}